### v1

# Adaptive Depth Networks with Skippable Sub-Paths

Abstract
Systematic adaptation of network depths at runtime can be an effective way to control inference latency and meet the resource condition of various devices. However, previous depth adaptive networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps. In this paper, we present an architectural pattern and training method for adaptive depth networks that can provide flexible accuracy-efficiency trade-offs in a single network. In our approach, every residual stage is divided into 2 consecutive sub-paths with different properties. While the first sub-path is mandatory for hierarchical feature learning, the other is optimized to incur minimal performance degradation even if it is skipped. Unlike previous adaptive networks, our approach does not iteratively self-distill a fixed set of sub-networks, resulting in significantly shorter training time. However, once deployed on devices, it can instantly construct sub-networks of varying depths to provide various accuracy-efficiency trade-offs in a single model. We provide a formal rationale for why the proposed architectural pattern and training method can reduce overall prediction errors while minimizing the impact of skipping selected sub-paths. We also demonstrate the generality and effectiveness of our approach with various residual networks, both from convolutional neural networks and vision transformers.
1 Introduction
Modern deep neural networks such as convolutional neural networks (CNNs) and transformers [44] provide state-of-the-art performance at high computational costs, and, hence, lots of efforts have been made to leverage those inference capabilities in various resource-constrained devices. Those efforts include compact architectures [21, 13], network pruning [14, 33], weight/activation quantization [26], knowledge distillation [19], to name a few. However, those approaches provide static accuracy-efficiency trade-offs, and, hence, it is infeasible to deploy one single model to meet devices with all kinds of resource-constraints.
There have been some attempts to provide predictable adaptability to neural networks by exploiting the redundancy in either network depths [24, 6], widths [59, 48], or both [47, 20]. However, one major difficulty with prior adaptive networks is that they are hard to train and require significantly longer training time than non-adaptive networks. For example, most adaptive networks select a fixed number of sub-networks of varying depths or width, and train them iteratively, mostly through self-distilling knowledge from the largest sub-network (also referred to as the super-net) [59, 48, 20, 43]. However, this iterative self-distillation takes long time and can generate conflicting training objectives for different parameter-sharing sub-networks, potentially resulting in worse performance [30, 10]. Further, unlike width-adaptation networks, no general principle has been proposed for how to select sub-networks for depth adaptation, since the effect of skipping individual layers has not been formally specified.
Figure 1: (a) At training time, the last half layers of every residual stage is optimized to have minimal performance degradation even if they are skipped. No explicit self-distillation is required for sub-networks, except the smallest base-net. (b) At test time, $2^{N_{r}}$ sub-networks become available in a single network by systematically skipping last half layers of $N_{r}$ residual stages. (c) Sub-networks instantly selected from a single network outperform counterpart individually trained networks due to regularization effect.
In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to residual networks, e.g., CNNs and transformers. In the proposed adaptive depth networks, every residual stage is divided into 2 consecutive sub-paths and they are trained to have different properties. While the first sub-paths are mandatory for hierarchical feature learning, the second sub-paths are optimized to incur minimal performance degradation even if they are skipped. More specifically, the second sub-path of every residual stage is optimized to preserve the feature distribution from its previous mandatory sub-path in order to minimize the performance degradation when it is skipped. During training, this property of the second sub-paths is enforced through skip-aware self-distillation, in which only one smallest sub-network, also referred as the base-net, is jointly trained in order to self-distill intermediate feature distributions at every residual stage, as shown in Figure-1-(a). This skip-aware self-distillation requires no explicit self-distillation or additional fine-tuning of individual sub-networks except the base-net, resulting in significantly shorter training time than previous adaptive networks. However, once trained, sub-networks with various depths can be selected instantly from a single network to meet the resource condition of devices, as shown in Figure 1-(b). Further, these sub-networks with various depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
In Section 3, we discuss the details of our architectural pattern and training algorithm, and show formally that the chosen sub-paths trained with our skip-aware self-distillation are optimized to reduce prediction errors while minimally changing the level of input features. In Section 4, we empirically demonstrate that our adaptive depth networks with skippable sub-paths outperform counterpart individual networks, both in CNNs and vision transformers, and achieve actual inference acceleration and energy-saving.
To authors’ best knowledge, this work is the first general approach to adaptive depth networks, providing a general principle for depth adaptation and a formal explanation on why layers can be skipped with minimal performance degradation.
Figure 2: Illustration of a residual stage with 2 sub-paths. While the first sub-path is mandatory for hierarchical feature learning, the second (yellow) sub-path can be skipped for efficiency. The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using skip-aware self-distillation. Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input to the next residual stage. In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped. This architectural pattern repeats in every residual stage throughout the network.
2 Related Work
Adaptive Networks: In most adaptive networks, parameter-sharing sub-networks are selected by adjusting either widths, depths, or resolutions [24, 59, 48, 22, 20, 30, 61, 47, 3]. For example, slimmable neural networks adjust channel widths of CNN models on the fly for accuracy-efficiency trade-offs and they exploit switchable batch normalization to handle multiple sub-networks [59, 58, 48]. Transformer-based adaptive depth networks have been proposed for language models to dynamically skip some of the layers during inference [20, 6]. However, in these adaptive networks, every target sub-network with varying widths or depths need to be trained explicitly, incurring significant training overheads and potential conflicts between sub-networks.
Dynamic networks [15] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaption of CNN models [51, 29, 12, 31, 53] and transformers [35, 55, 8, 18]. In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layers (or blocks) that determine if the layers can be skipped [9, 51, 45, 50]. These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs. However, the learned policy for skipping layers is opaque to users and does not provide a formal description of when and which layers can be skipped for a given input. Therefore, the network depths cannot be adapted in a predictable manner to meet the resource condition of target devices.
Residual Blocks with Shortcuts: Since the introduction of ResNets [16], residual blocks with shortcuts have received extensive attention because of their ability to train very deep networks, and have been chosen by many CNNs [38, 41] and transformers [44, 5, 34]. In [46], Veit et al. argue that identity shortcuts make exponential paths and results in an ensemble of shallower sub-networks. This thought is supported by the fact that removing individual residual blocks at test time does not significantly affect performance, and it has been further exploited to train deep networks [23, 52]. Other works argue that identity shortcuts enable residual blocks to perform iterative feature refinement, where each block improves slightly but keeps the semantic of the representation of the previous layer [11, 27]. Our work build upon those views on residual blocks with shortcuts and further extend them for adaptive depth networks by introducing an architectural pattern and training method that exploits the properties of residual blocks more explicitly for selected sub-paths.
3 Adaptive Depth Networks
We first present an architectural pattern and training details to build adaptive depth networks. And, then, we discuss the theoretic rationale for how depth adaption can be achieved with minimal performance degradation.
3.1 Architectural Pattern for Depth Adaptation
In typical hierarchical residual networks such as ResNets [16] and Swin transformers [34], the $s$-th residual stage is consisted of $L$ identical residual blocks, which transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
$\underbrace{\underbrace{\mathbf{h}^{s}_{1}+\underbrace{F_{1}(\mathbf{h}^{s}_{1})+...+{F}^{s}_{L/2}(\mathbf{h}^{s}_{L/2})}_{\mathbf{F}^{s}_{base}}}_{=\mathbf{h}^{s}_{base}}+\underbrace{...+F_{L}(\mathbf{h}^{s}_{L})}_{\mathbf{F}^{s}_{skippable}}}_{=\mathbf{h}^{s}_{super}}=\mathbf{h}^{s}$ (1)
While a block with a residual function $F_{\ell}$ ($\ell=1,...,L$) learns hierarchical features as traditional compositional networks [40], previous literature [27, 11] demonstrates that a residual function also tend to learn a function that refines already learned features at the same feature level. If a residual block mostly performs feature refinement while not changing the level of input features, the performance of the residual network is not significantly affected by dropping the block at test time [23, 52]. However, in typical residual networks, most residual blocks tend to refine features while learning new level features as well, and, hence, random dropping of residual blocks at test time degrades the performance significantly. Therefore, we hypothesize that if some designated residual blocks can be encouraged during training to focus more on feature refinement, then these blocks can be skipped to save computation at marginal loss of prediction accuracy at test time.
To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two consecutive sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2. While $\mathbf{F}^{s}_{base}$ learns feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}$. Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference. However, since layers in $\mathbf{F}^{s}_{skippable}$ only refine $\mathbf{h}^{s}_{base}$, they can be skipped to save computation. If $\mathbf{F}^{s}_{skippable}$ is skipped, then intermediate features $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage. Therefore, the overall network depth can be adjusted by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{s})$ or not.
In Section 4.1, we show that this architectural pattern for adaptive depth networks is generally applicable to a wide range of residual networks.
3.2 Skip-Aware Self-Distillation
Preserving the feature level of $\mathbf{h}^{s}_{base}$ in $\mathbf{F}^{s}_{skippable}$ implies, more specifically, that two feature representations $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$ have similar distributions over training input $\mathbf{X}$. If $\mathbf{h}^{s}_{super}$ and $\mathbf{h}^{s}_{base}$ have similar distributions, skipping the sub-path $\mathbf{F}^{s}_{skippable}$ during inference results in minimal internal covariate shifts [25] to the following network layers. Kullback-Leibler (KL) divergence measures how different two distributions are over the same random variable, and, hence, we use Equation 2 to measure the similarity of two distributions, $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$, over input $\mathbf{X}$:
$\underset{x\in\mathbf{X}}{D_{KL}}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ (2)
Algorithm 1 shows our training method, called skip-aware self-distillation, in which Equation 2 is included in the loss function while the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are jointly trained  [^1] Some frameworks, such as Pytorch’s distributed data parallel (DDP), do not support two consecutive forwards passes followed by a single backward pass: $super.forward\rightarrow base.fowrard\rightarrow loss.backward$. For such frameworks, we adapt the algorithm to $super.forward\rightarrow loss_{super}.backward\rightarrow base.forward\rightarrow loss_{sub\_path}.backward$. The latter method has similar results and is more memory-efficient. . In Algorithm 1, the forward function of the proposed adaptive depth networks $\mathbf{M}$ accepts an extra argument, ’skip’, that controls which residual stages skip their skippable sub-paths. For example, if $\mathbf{M}$ has 4 residual stages, its base-net is selected by passing ‘skip=[True, True, True, True]’.
Algorithm 1 Training algorithm for an adaptive depth network $\mathbf{M}$. Two parameter-sharing sub-networks, or super-net and base-net, are jointly trained using skip-aware self-distillation. $N_{r}$ denotes the number of residual stages.
1:Initialize an adaptive depth network $\mathbf{M}$
2:for $i=1$ to $n_{iters}$ do
3:     Get next mini-batch of data $\mathbf{x}$ and label $\mathbf{y}$
4:     $optimizer.zero\_grad()$
[…]
6:     $loss_{super}=criterion(\mathbf{y},\mathbf{\hat{y}}_{super})$
[…]
8:     $loss_{base}=\sum_{s=1}^{N_{r}}D_{KL}(\mathbf{h}^{s}_{super}\|\mathbf{h}^{s}_{base})\newline \hskip 50.00008pt+D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$
9:     $loss=\alpha loss_{super}+(1-\alpha)loss_{base}$
10:     $loss.backward()$
11:     $optimizer.step()$
12:end for
In steps 5 and 7, the forward passes of the super-net and the base-net are executed for the same input $\mathbf{x}$. Intermediate features $\mathbf{h_{super}}$ and $\mathbf{h_{base}}$ are also obtained during the forward passes of $\mathbf{M}$. In step 8, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in Equation 2 is included in the loss function $loss_{base}$. By minimizing $loss_{base}$, two feature representations $\mathbf{h}^{s}_{super}$ and $\mathbf{h}^{s}_{base}$, are explicitly enforced to have similar distributions for the same input $\mathbf{x}$. Further, this step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage. Therefore, $\mathbf{h}^{s}_{base}$ is expected to learn more compact representation from $\mathbf{h}^{s}_{super}$. The hyperparameter $\alpha$ controls the strength of this skip-aware self-distillation. In step 8, $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is included in the loss function for further distillation effect from the super-net to the sub-networks. Due to the architectural pattern of interleaving the mandatory and the skippable sub-paths, minimizing $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ also minimizes $D_{KL}(\mathbf{h}_{super}\|\mathbf{h}_{base})$ implicitly. Experimental result shows that similar results can be achieved when only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is used in $loss_{base}$ (Section 4.5). This implicit approach is useful when the extraction of intermediate features is tricky.
In Algorithm 1, only two sub-networks are involved for the training of $\mathbf{M}$, and, hence, the total training time is no greater than training two sub-networks individually. However, at test time, sub-networks with various depths can be selected on the fly by systematically skipping sub-paths in residual stages. For example, in a network with 4 residual stages, $16\ (=2^{4})$ parameter-sharing sub-networks can be selected by varying the skip argument. In contrast, prior adaptive networks supporting 16 parameter-sharing sub-networks need to perform 16 explicit self-distillation from the super-net to the sub-networks [48, 59, 20]. We demonstrate this result in Section 4.3.
3.3 Formal Analysis of Skippable Sub-Paths
$D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in Equation 2 can be trivially minimized if residual blocks in $\mathbf{F}^{s}_{skippable}$ learn identity functions, or $\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})=\mathbf{h}^{s}_{base}$. However, since the super-net is jointly trained with the loss function $loss_{super}$, the residual functions in $\mathbf{F}^{s}_{skippable}$ cannot simply be an identity function. Then, what do the residual functions in $\mathbf{F}^{s}_{skippable}$ learn during training? This can be further investigated through Taylor expansion [27]. For our adaptive depth networks, a loss function $\mathcal{L}$ used for training the super-net can be approximated with Taylor expansion as follows:
$\mathcal{L}(\mathbf{h}^{s}_{super})=\mathcal{L}\{\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})\}$ (3)
=$=\mathcal{L}\{\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})+F_{L}(\mathbf{h}^{s}_{L})\}$ (4)
$\approx\mathcal{L}\{\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +F_{L}(\mathbf{h}^{s}_{L})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{L})}{\partial\mathbf{h}^{s}_{L}}+\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ (5)
In Equation 5, the loss function is expanded around $\mathbf{h}^{s}_{L}$, or $\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})$. Only the first order term is left and all high order terms, such as $F_{L}(\mathbf{h}_{L}^{s})^{2}\cdot\frac{{\partial^{2}\mathcal{L}(\mathbf{h}^{s}_{L})}}{2{\partial(\mathbf{h}^{s}_{L})^{2}}}$, are absorbed in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$.
The terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude. In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude. In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the skip-aware self-distillation, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation. If we similarly keep expanding the loss function around $\mathbf{h}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
$\mathcal{L}(\mathbf{h}^{s}_{super})\approx\mathcal{L}(\mathbf{h}^{s}_{base})+\sum_{j={L/2+1}}^{L}F_{j}(\mathbf{h}^{s}_{j})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{j})}{\partial\mathbf{h}^{s}_{j}}$ (6)
[…]
$F_{j}(\mathbf{h}^{s}_{j})\simeq\mathbf{-}\textrm{ }\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{j})}{\partial\mathbf{h}^{s}_{j}}\textrm{ }(j=L/2+1,...,L)$ (7)
In other words, the residual functions in the skippable sub-paths reduce the loss $\mathcal{L}(\mathbf{h}^{s}_{base})$ iteratively during inference while preserving the feature distribution of $\mathbf{h}^{s}_{base}$.
Considering this result, we can conjecture that, with our architectural pattern and training method, layers in $\mathbf{F}^{s}_{skippable}$ learn functions that refine input features $\mathbf{h}^{s}_{base}$ iteratively for better inference accuracy while minimally changing the distribution of $\mathbf{h}^{s}_{base}$. Therefore, skipping $\mathbf{F}^{s}_{skippable}$ only slightly reduces prediction accuracy.
3.4 Skip-aware Batch Normalization
Originally, batch normalization (BN) [25] was proposed to handle internal covariate shift during training non-adaptive networks by normalizing features. In our adaptive depth networks, however, internal covariate shifts can occur during inference in mandatory sub-paths if different sub-networks are selected. To handle potential internal covariate shifts, switchable BN operators, called skip-aware BNs, are used in mandatory sub-paths. For example, at each residual stage, two sets of BNs are available for the mandatory sub-path, and they are switched depending on whether its skippable sub-path is skipped or not.
The effectiveness of switchable BNs has been demonstrated in networks with adaptive widths [57, 48] and adaptive resolutions [62]. However, in previous adaptive networks, $N$ sets of switchable BNs are required in every layer to support $N$ parameter-sharing sub-networks. Such a large number of switchable BNs not only requires more parameters, but also makes the training process complicated since $N$ sets of switchable BNs need to be trained iteratively during training. In contrast, in our adaptive depth networks, every mandatory sub-path needs only two sets of switchable BNs, regardless of the number of supported sub-networks. This reduced number of switchable BNs significantly simplifies the training process as shown in Algorithm 1. Furthermore, the amount of parameters for skip-aware BNs is negligible. For instance, in ResNet50, skip-aware BNs increase the parameters by 0.07%.
Transformers [44, 34] exploit layer normalization (LN) instead of BNs and naive replacement of LNs to BNs incurs instability during training [54]. Therefore, for our adaptive depth transformers, we apply switchable LN operators in mandatory sub-paths instead of switchable BNs.
4 Experiments
To demonstrate the generality and effectiveness of our approach for adaptive depth networks, we conduct experiments on various networks and vision tasks.
4.1 Networks
We use four representative residual networks as base models to apply the proposed architecture pattern in Section 3.1: MobileNet V2 [38] is a lightweight CNN model, ResNet [16] is a larger CNN model, and ViT [5] and Swin-T [34] are representative vision transformers. All base models except ViT have 4 residual stages, each with 2 $\sim$ 6 (residual or encoder) blocks. So, according to the proposed architectural pattern, every residual stage is evenly divided into 2 sub-paths for depth adaptation. Since ViT does not define residual stages, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and desiginate the last encoder block of each group as a skippable sub-path.
We use the suffix ‘-ADN’ to denote our adaptive depth networks. Since our adaptive depth networks have many parameter-sharing sub-networks in a single network, we indicate which sub-network is used for evaluation in parenthesis, e.g., ResNet50-ADN (super-net). For our sub-networks, boolean values are also used to indicate in which residual stages the sub-paths are skipped. For example, ResNet50-ADN (base-net) is equivalent to ResNet50-ADN (TTTT).
Table 1: Results on ImageNet. Individual networks with the suffix ‘-Base’ have the same depths as the base-nets of corresponding adaptive depth networks. Latency is measured on an RTX 3090 (batch size = 64).
Model Params FLOPs Acc@1 Latency
(M) (G) (%) (msec)
ResNet50-ADN (super-net) 25.58 4.11 77.6 43.2
ResNet50-ADN (base-net) 2.58 76.1 27.6
ResNet50 (individual) 25.56 4.11 76.7 43.2
ResNet50-Base (individual) 17.11 2.58 75.0 27.6
MV2-ADN (super-net) 3.72 0.32 72.7 15.5
MV2-ADN (base-net) 0.25 70.7 13.2
MV2 (individual) 3.50 0.32 72.1 15.5
MV2-Base (individual) 2.99 0.25 70.7 13.2
ViT-b/16-ADN (super-net) 86.59 17.58 81.3 138.6
ViT-b/16-ADN (base-net) 11.76 79.4 90.7
ViT-b/16 (individual) 86.57 17.58 81.1 138.6
ViT-b/16-Base (individual) 67.70 11.76 78.7 90.7
Swin-T-ADN (super-net) 28.30 4.49 81.8 74.4
Swin-T-ADN (base-net) 2.69 78.7 39.9
Swin-T (individual) 28.29 4.49 81.5 74.4
Swin-T-Base (individual) 15.63 2.69 78.6 39.9
4.2 ImageNet Classification
We evaluate our method on ILSVRC2012 dataset [37] that has 1000 classes. The dataset consists of 1.28M training and 50K validation images. For CNN models, we follow most training settings in the original papers [16, 38], except that ResNet models are trained for 150 epochs. ViT and Swin-T are trained for 300 epochs, following DeiT’s training recipe [42, 1]. However, in Swin-T-ADN, we disable stochastic depths [23] for the base-net since the strategy of random dropping of residual blocks conflicts with our approach to skipping sub-paths. For fair comparison, our adaptive depth networks and corresponding individual networks are trained in the same training settings. The hyperparameter $\alpha$ in Algorithm 1 is set to $0.5$ for all networks.
The results in Table 1 show that our adaptive depth networks outperform individual counterpart networks even though our sub-networks share parameters in a single network. Further, our results with vision transformers demonstrate that our approach is generally applicable to residual networks and compatible with their state-of-the-art training techniques such as DeiT’s training recipe [42]. We conjecture that this performance improvement results from effective distillation of knowledge from $\textbf{h}^{s}_{super}$ to $\textbf{h}^{s}_{base}$ at each residual stage and the iterative feature refinement at skippable sub-paths, shown in Equation 7.
In Table 2 and Figure 3, several state-of-the-art efficient inference methods and dynamic networks are compared with our sub-networks of adaptive depth networks. The result demonstrates that our adaptive depth networks match or outperform many state-of-the-art static and dynamic networks across varying depth ranges.
In Figure 3, sub-networks of ResNet50-ADN are compared to the counterpart individual ResNets of equivalent depths. In particular, it should be noted that ResNets trained with knowledge distillation in the same training settings has worse performance than individual ResNets trained without knowledge distillation. As reported in previous works, successful knowledge distillation requires a patient and long training [2], and straightforward knowledge distillation using ImageNet does not improve the performance of student models [60, 4]. In contrast, our ResNet50-ADN trained with skip-aware self-distillation achieves better performance than counterpart ResNets. This demonstrates that the high performance of adaptive depth networks does not simply come from the distillation effect, but from the effective combination of the proposed architectural pattern and the skip-aware self-distillation strategy.
Figure 3: Comparison with various approaches on ImageNet. Sub-networks of our adaptive depth networks match or outperform other approaches on ImageNet. ResNet50 (individual) shows individual networks that have equivalent network depths as the sub-networks of ResNet50-ADN. ResNet50 (KD individual) shows individual networks trained using knowledge distillation from pretrained Pytorch ResNet50. In legend, $\dagger$ denotes static compression methods, and $\star$ denotes input-dependent dynamic networks.
Table 2: Our base-nets are compared with state-of-the-art efficient inference methods for various networks. $\dagger$ denotes static compression methods, $\ast$ denotes width-adaptation networks, and $\star$ denotes input-dependent dynamic networks. Note that many of these state-of-the-art methods exploit various non-canonical training techniques, such as iterative retraining and post-processing, while our base-nets are selected from adaptive depth networks without any fine-tuning.
[…]
GReg-1 [49$\dagger$ 2.8G 33% 76.1%
DR-ResNet50 ($\alpha$=2.0) [62$\star$ 2.3G 44% 75.3%
DepGraph [7$\dagger$ 2.0G 51% 75.8%
ResNet50-ADN (base-net) 2.6G 37% 76.1%
AlphaNet-0.75x [48] $\ast$ 0.21G 34% 70.5%
MV2-ADN (base-net) 0.25G 22% 70.7%
Ada-DeiT-S [35$\star$ 2.3G - 77.3%
WDPruning-0.3-11 [56$\dagger$ 2.6G - 78.3%
DeiT-S+ATS (not fine-tuned) [8$\star$ 3.0G - 78.5%
SPViT-Swin-Ti [28$\star$ 3.3G 27% 80.1%
Swin-T-ADN (base-net) 2.7G 40% 78.7%
4.3 Training Time
Table 3: Training time (1 epoch), measured on a Nvidia RTX 4090 (batch size = 128). All compared models are based on ResNet50 or have similar FLOPs and depths.
Model Training time (min) # of sub-networks
ResNet50 32.9 1
ResNet50-Base 22.7 1
ResNet50-ADN (ours) 54.7 $2^{4}$ (varying depths)
MSDNet [24] 67.1 9 (early-exit branches)
S-ResNet50 [59] 82.5 4 (varying widths)
SkipNet [50] 90.6 dynamic depths
One important advantage of our approach is that our adaptive depth networks require significantly shorter training time than other adaptive networks. Table 3 shows that training our ResNet50-ADN takes similar time to train two individual networks combined. In contrast, the compared adaptive networks requires much longer training time than ours. MSDNet [24] represents adaptive depth networks with multiple early-exiting branches and classifiers. SkipNet [50] represents input-dependent dynamic networks that drop layers for ‘easy’ inputs. In S-ResNet50 [59], 4 sub-networks with varying widths are trained through iterative self-distillation from its super-net.
Despite significantly shorter training time is required, our adaptive depth networks can support various sub-networks at test time. Figure 4-(left) shows the performance of ResNet50-ADN when its depth is varied at test time. Among them, only the super-net and the base-net are explicitly trained in Algorithm 1. Other sub-networks are selected on the fly at test time by skipping sub-paths in varying number of residual stages. Although these sub-networks are not trained explicitly, they show graceful degradation of performance as the depth of sub-networks becomes gradually shallower. However, although ResNet50-ADN support $2^{4}$ sub-networks in theory, Figure 4-(right) shows that some sub-networks are more useful than the others. For instance, even though ResNet50-ADN(TFFF) and ResNet50-ADN(FFFT) have the same network depths, ResNet50-ADN(TFFF) has about $0.58\%$ higher top-1 accuracy. This result shows that skipping sub-paths in later stages is more detrimental to the performance in general.
Figure 4: Validation accuracy of our sub-networks of varying depths. The boolean values below the bars indicate the residual stages where the sub-paths are skipped.
4.4 Performance on Devices
Figure 5 shows the performance of adaptive networks in an actual device. The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50 [59], a representative width-adaptation network. The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption. For example, in ResNet50-ADN, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%. In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
Figure 5: Inference latency and energy consumption of adaptive networks, measured on Nvidia Jetson Orin Nano (batch size = 1). Although adjusting the network width can support a much wider FLOPs adaptation range, depth adaptation of ResNet50-ADN is more effective for actual inference acceleration and saving energy.
Table 4: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. In the last row, only $D_{KL}$ with softmax outputs is used for the skip-aware self-distillation. $\downarrow$ and $\uparrow$ in parentheses are comparisons to individual networks.
Skip-aware Skip-aware ResNet50 Acc@1 ViT-b/32 Acc@1
self-distillation BNs / LNs Super-net Base-net Super-net Base-net
75.2% ($\downarrow$ 1.5%) 72.2% ($\downarrow$ 2.8%) 75.7% ($\downarrow$0.2%) 74.1% ($\uparrow$ 0.3%)
O 76.1% ($\downarrow$ 0.6%) 74.9 % ($\downarrow$ 0.1%) 76.4% ($\uparrow$0.5%) 74.3% ($\uparrow$0.5%)
O 76.6% ($\downarrow$ 0.1%) 75.1% ($\uparrow$ 0.1%) 76.0% ($\uparrow$0.1%) 74.3% ($\uparrow$0.5%)
O O 77.6% ($\uparrow$ 0.9%) 76.1% ($\uparrow$ 1.1%) 76.6% ($\uparrow$ 0.8%) 74.3% ($\uparrow$0.5%)
$\triangle$ O 77.6% ($\uparrow$ 0.9%) 75.9% ($\uparrow$ 0.9%)
4.5 Ablation Study
We conduct ablation study on ImageNet classification to investigate the influence of two key components of the proposed adaptive depth networks: (1) skip-aware self-distillation and (2) skip-aware BN/LNs. When our skip-aware self-distillation is not applied, the loss in Algorithm 1 is modified to $loss=\frac{1}{2}\{criterion(\mathbf{y},\mathbf{\hat{y}}_{super})+criterion(\mathbf{y},\mathbf{\hat{y}}_{base})\}$ for joint training of the super-net and the base-net. Table 4 shows the results. For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than individual networks by 1.5% and 2.8%, respectively. This result shows the difficulty of joint training sub-networks for adaptive networks. When one of the two components is applied individually, the performance is still slightly worse than individual networks’. In the third row, when both skip-aware self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net. Finally, the last row shows the result when only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is used for skip-aware self-distillation. Without explicit distillation of intermediate features, slightly lower performance is observed in the base-net. The results with ViT-b/32 show that switchable layer normalization has similar effect in vision transformers.
(a) Original ResNet50
(b) ResNet50-ADN (super-net)
(c) ResNet50-ADN (base-net)
Figure 6: Class Activation Maps of the 3rd residual stages of ResNet50s. (a) Original ResNet50’s activation regions change gradually across all blocks. (b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features. In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target. (c) The base-net’s final activation map is very similar to super-net’s final activation map, implying that they are at the same feature level.
4.6 Visual Analysis of Sub-Paths
To investigate how our training method affects feature representations in the mandatory and the skippable sub-paths, we visualize the activation of 3rd residual stage of ResNet50-ADN using Grad-CAM [39]. The 3rd residual stage of ResNet50-ADN has 6 residual blocks and the last three blocks are skippable. In Figure 6-(a), the activation regions of original ResNet50 changes gradually across all consecutive blocks. In contrast, in Figure 6-(b), ResNet50-ADN (super-net) manifests very different activation regions in two sub-paths. In the first three residual blocks, we can observe lots of hot activation regions in wide areas, suggesting active learning of new level features. In contrast, significantly less activation regions are found in the skippable last three blocks and they are gradually concentrated around the target object, demonstrating the refinement of learned features. Further, in Figure 6-(c), we can observe that the final activation map of the ResNet50 (base-net) is very similar to the super-net’s final activation map in Figure 6-(b). This implies that they have similar distributions for the same inputs, as suggested in Section 3.3.
4.7 Object Detection and Instance Segmentation
Table 5: Object detection and instance segmentation results on MS COCO dataset.
Detector Backbone FLOPs Individual Networks ResNet50-ADN (ours)
Box AP Mask AP Box AP Mask AP
Faster-RCNN ResNet50 207.07G 36.4 37.8
[36] ResNet50-Base 175.66G 32.4 34.0
Mask-RCNN ResNet50 260.14G 37.2 34.1 38.3 34.1
[17] ResNet50-Base 228.73G 32.7 29.9 34.1 31.2
In order to investigate the generalization ability of our approach, we use MS COCO 2017 datasets on object detection and instance segmentation tasks using representative detectors. We compare individual ResNet50 and our adaptive depth ResNet50-ADN as backbone networks of the detectors. For training of detectors, we use Algorithm 1 with slight adaptation. For object detection, the intermediate features $\textbf{h}_{base}^{s}$ and $\textbf{h}_{super}^{s}(s=1..N_{r})$ can be obtained directly from backbone network’s feature pyramid networks (FPN) [32], and, hence, a wrapper function is not required to extract intermediate features. All networks are trained on train2017 for 12 epochs from ImageNet pretrained weights, following the training settings suggested in [32]. Table 5 shows the results on val2017 containing 5000 images. Our adaptive depth backbone networks still outperform individual static backbone networks in terms of COCO’s standard metric AP.
5 Conclusions
We propose an architectural pattern for adaptive depth networks and its training method that renders selected sub-paths of the network to incur minimal performance degradation even if they are skipped during inference. Our approach does not train a fixed set of sub-networks iteratively, and, hence significantly shorter training time is required than previous adaptive networks. However, once deployed on devices, it can instantly construct sub-networks of various depths to provide various accuracy-efficiency trade-offs in a single network. We show formally that the proposed architectural pattern and training method reduces overall prediction errors while minimizing the impact of skipping sub-paths. We also empirically demonstrate the generality and effectiveness of our approach using representative convolutional neural networks and vision transformers.

### v2

@@
 Abstract
-Systematic adaptation of network depths at runtime can be an effective way to control inference latency and meet the resource condition of various devices.
-However, previous depth adaptive networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps.
-In this paper, we present an architectural pattern and training method for adaptive depth networks that can provide flexible accuracy-efficiency trade-offs in a single network.
-In our approach, every residual stage is divided into 2 consecutive sub-paths with different properties.
-While the first sub-path is mandatory for hierarchical feature learning, the other is optimized to incur minimal performance degradation even if it is skipped.
-Unlike previous adaptive networks, our approach does not iteratively self-distill a fixed set of sub-networks, resulting in significantly shorter training time.
-However, once deployed on devices, it can instantly construct sub-networks of varying depths to provide various accuracy-efficiency trade-offs in a single model.
-We provide a formal rationale for why the proposed architectural pattern and training method can reduce overall prediction errors while minimizing the impact of skipping selected sub-paths.
-We also demonstrate the generality and effectiveness of our approach with various residual networks, both from convolutional neural networks and vision transformers.
+Predictable adaptation of network depths can be an effective way to control inference latency and meet the resource condition of various devices.
+However, previous adaptive depth networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps.
+In this paper, we present a practical approach to adaptive depth networks that is applicable to various networks with minimal training effort.
+In our approach, every hierarchical residual stage is divided into two sub-paths, and they are trained to acquire different properties through a simple self-distillation strategy.
+While the first sub-path is essential for hierarchical feature learning, the second one is trained to refine the learned features and minimize performance degradation if it is skipped.
+Unlike prior adaptive networks, our approach does not train every target sub-network in an iterative manner.
+At test time, however, we can connect these sub-paths in a combinatorial manner to select sub-networks of various accuracy-efficiency trade-offs from a single network.
+We provide a formal rationale for why the proposed training method can reduce overall prediction errors while minimizing the impact of skipping sub-paths.
+We demonstrate the generality and effectiveness of our approach with convolutional neural networks and transformers.
 1 Introduction
@@
 For example, most adaptive networks select a fixed number of sub-networks of varying depths or width, and train them iteratively, mostly through self-distilling knowledge from the largest sub-network (also referred to as the super-net) [10, 11, 13, 14].
-However, this iterative self-distillation takes long time and can generate conflicting training objectives for different parameter-sharing sub-networks, potentially resulting in worse performance [30, 10].
-Further, unlike width-adaptation networks, no general principle has been proposed for how to select sub-networks for depth adaptation, since the effect of skipping individual layers has not been formally specified.
-Figure 1: (a) At training time, the last half layers of every residual stage is optimized to have minimal performance degradation even if they are skipped.
-No explicit self-distillation is required for sub-networks, except the smallest base-net.
-(b) At test time, $2^{N_{r}}$ sub-networks become available in a single network by systematically skipping last half layers of $N_{r}$ residual stages.
-(c) Sub-networks instantly selected from a single network outperform counterpart individually trained networks due to regularization effect.
-In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to residual networks, e.g., CNNs and transformers.
-In the proposed adaptive depth networks, every residual stage is divided into 2 consecutive sub-paths and they are trained to have different properties.
+However, this exhaustive self-distillation takes long time and can generate conflicting training objectives for different parameter-sharing sub-networks, potentially resulting in worse performance [15, 16].
+Moreover, unlike width-adaptation networks, there is no established principle for adapting depths since the impact of skipping individual layers has not been formally defined.
+Figure 1: (a) During training, each residual stage of a network is divided into two sub-paths.
+The layers in the second sub-paths are optimized to minimize performance degradation even if they are skipped.
+(b) At test time, we can instantly select $2^{N_{r}}$ parameter-sharing sub-networks from a single network by systematically skipping the last half layers of $N_{r}$ residual stages.
+(c) The sub-networks selected from a single network form a better Pareto frontier compared to individual networks.
+In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to various networks, e.g., CNNs and transformers.
+In the proposed adaptive depth networks, every residual stage is divided into two sub-paths and the sub-paths are trained to have different properties.
 While the first sub-paths are mandatory for hierarchical feature learning, the second sub-paths are optimized to incur minimal performance degradation even if they are skipped.
-More specifically, the second sub-path of every residual stage is optimized to preserve the feature distribution from its previous mandatory sub-path in order to minimize the performance degradation when it is skipped.
-During training, this property of the second sub-paths is enforced through skip-aware self-distillation, in which only one smallest sub-network, also referred as the base-net, is jointly trained in order to self-distill intermediate feature distributions at every residual stage, as shown in Figure-1-(a).
-This skip-aware self-distillation requires no explicit self-distillation or additional fine-tuning of individual sub-networks except the base-net, resulting in significantly shorter training time than previous adaptive networks.
-However, once trained, sub-networks with various depths can be selected instantly from a single network to meet the resource condition of devices, as shown in Figure 1-(b).
-Further, these sub-networks with various depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
-In Section 3, we discuss the details of our architectural pattern and training algorithm, and show formally that the chosen sub-paths trained with our skip-aware self-distillation are optimized to reduce prediction errors while minimally changing the level of input features.
-In Section 4, we empirically demonstrate that our adaptive depth networks with skippable sub-paths outperform counterpart individual networks, both in CNNs and vision transformers, and achieve actual inference acceleration and energy-saving.
-To authors’ best knowledge, this work is the first general approach to adaptive depth networks, providing a general principle for depth adaptation and a formal explanation on why layers can be skipped with minimal performance degradation.
-Figure 2: Illustration of a residual stage with 2 sub-paths.
-While the first sub-path is mandatory for hierarchical feature learning, the second (yellow) sub-path can be skipped for efficiency.
-The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using skip-aware self-distillation.
-Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input to the next residual stage.
-In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped.
-This architectural pattern repeats in every residual stage throughout the network.
+In order to enforce this property of the second sub-paths, we propose a simple self-distillation strategy, in which only the largest sub-network (or, super-net) and the smallest sub-network (or, base-net) are exclusive used as a teacher and a student, respectively, as shown in Figure-1-(a).
+The proposed self-distillation strategy does not require exhaustive training of every target sub-network, resulting in significantly shorter training time than prior adaptive networks.
+Nevertheless, at test time, sub-networks with various depths can be selected instantly from a single network by connecting these sub-paths in a combinatorial manner, as shown in Figure 1-(b).
+Further, these sub-networks with varying depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
+In Section 3, we discuss the details of our architectural pattern and training algorithm, and show formally that the selected sub-paths trained with our self-distillation strategy are optimized to reduce prediction errors while minimally changing the level of input features.
+In Section 4, we empirically demonstrate that our adaptive depth networks outperform counterpart individual networks, both in CNNs and vision transformers, and achieve actual inference acceleration and energy-saving.
+To authors’ best knowledge, this is the first approach to adaptive networks that provides a general principle and theoretic basis for how depth adaptation can be supported with minimal performance degradation and training effort.
+This allows us to avoid typical exhaustive training of target sub-networks and instead construct sub-networks of varying depths from specifically trained sub-paths.
+We anticipate these advances will facilitate future research and applications of adaptive networks.
 2 Related Work
-Adaptive Networks: In most adaptive networks, parameter-sharing sub-networks are selected by adjusting either widths, depths, or resolutions [24, 59, 48, 22, 20, 30, 61, 47, 3].
+Adaptive Networks: In most adaptive networks, parameter-sharing sub-networks are instantly selected by adjusting either widths, depths, or resolutions [8, 10, 11, 17, 13, 15, 18, 12, 19].
 For example, slimmable neural networks adjust channel widths of CNN models on the fly for accuracy-efficiency trade-offs and they exploit switchable batch normalization to handle multiple sub-networks [10, 20, 11].
@@
 However, in these adaptive networks, every target sub-network with varying widths or depths need to be trained explicitly, incurring significant training overheads and potential conflicts between sub-networks.
-Dynamic networks [15] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaption of CNN models [51, 29, 12, 31, 53] and transformers [35, 55, 8, 18].
+Dynamic Networks: Dynamic networks [21] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaptation of CNN models [22, 23, 24, 25, 26] and transformers [27, 28, 29, 30].
 In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layers (or blocks) that determine if the layers can be skipped [31, 22, 32, 33].
 These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs.
-However, the learned policy for skipping layers is opaque to users and does not provide a formal description of when and which layers can be skipped for a given input.
+However, the learned policy for skipping layers is opaque to users and does not provide a formal description of which layers can be skipped for a given input.
 Therefore, the network depths cannot be adapted in a predictable manner to meet the resource condition of target devices.
-Residual Blocks with Shortcuts: Since the introduction of ResNets [16], residual blocks with shortcuts have received extensive attention because of their ability to train very deep networks, and have been chosen by many CNNs [38, 41] and transformers [44, 5, 34].
-In [46], Veit et al.
-argue that identity shortcuts make exponential paths and results in an ensemble of shallower sub-networks.
+Residual Blocks with Shortcuts: Since the introduction of ResNets [34], residual blocks with identity shortcuts have received extensive attention because of their ability to train very deep networks, and have been chosen by many CNNs [35, 36] and transformers [1, 37, 38].
+Veit et al.
+[39] argue that identity shortcuts make exponential paths and results in an ensemble of shallower sub-networks.
 This thought is supported by the fact that removing individual residual blocks at test time does not significantly affect performance, and it has been further exploited to train deep networks [40, 41].
 Other works argue that identity shortcuts enable residual blocks to perform iterative feature refinement, where each block improves slightly but keeps the semantic of the representation of the previous layer [42, 43].
-Our work build upon those views on residual blocks with shortcuts and further extend them for adaptive depth networks by introducing an architectural pattern and training method that exploits the properties of residual blocks more explicitly for selected sub-paths.
+Our work build upon those views on residual blocks with shortcuts and further extend them for adaptive depth networks by introducing a training method that enforces the properties of residual blocks more explicitly for skippable sub-paths.
 3 Adaptive Depth Networks
-We first present an architectural pattern and training details to build adaptive depth networks.
-And, then, we discuss the theoretic rationale for how depth adaption can be achieved with minimal performance degradation.
+We first present the architecture and training details of adaptive depth networks.
+Then, we discuss the theoretic rationale for how depth adaptation can be achieved with minimal performance degradation.
+Figure 2: Illustration of a residual stage with two sub-paths.
+While the first (blue) sub-path is mandatory for hierarchical feature learning, the second (orange) sub-path can be skipped for efficiency.
+The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using the proposed self-distillation strategy.
+Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input $\mathbf{h}^{s}$ to the next residual stage.
+In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped.
+This architectural pattern is applicable to both CNNs and transformers.
 3.1 Architectural Pattern for Depth Adaptation
-In typical hierarchical residual networks such as ResNets [16] and Swin transformers [34], the $s$-th residual stage is consisted of $L$ identical residual blocks, which transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
-$\underbrace{\underbrace{\mathbf{h}^{s}_{1}+\underbrace{F_{1}(\mathbf{h}^{s}_{1})+...+{F}^{s}_{L/2}(\mathbf{h}^{s}_{L/2})}_{\mathbf{F}^{s}_{base}}}_{=\mathbf{h}^{s}_{base}}+\underbrace{...+F_{L}(\mathbf{h}^{s}_{L})}_{\mathbf{F}^{s}_{skippable}}}_{=\mathbf{h}^{s}_{super}}=\mathbf{h}^{s}$ (1)
+In hierarchical networks, such as ResNets [34] and Swin transformers [38], there are typically 4 to 7 consecutive residual stages.
+In a network with $N_{r}$ residual stages, the $s$-th ($s=1,...,N_{r}$) residual stage consists of $L$ identical residual blocks  [^1] In transformers, residual blocks are called encoder-decoder blocks.
+In most networks, $L$ is usually 3 to 6.
+that transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
+$\underbrace{\mathbf{h}^{s}_{1}}_{=\mathbf{h}^{s-1}}+\underbrace{F_{1}(\mathbf{h}^{s}_{1})+...+{F}^{s}_{L/2}(\mathbf{h}^{s}_{L/2})}_{\mathbf{F}^{s}_{base}}+\underbrace{F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})...+F_{L}(\mathbf{h}^{s}_{L})}_{\mathbf{F}^{s}_{skippable}}=\mathbf{h}^{s}$ (1)
 While a block with a residual function $F_{\ell}$ ($\ell=1,...,L$) learns hierarchical features as traditional compositional networks [44], previous literature [43, 42] demonstrates that a residual function also tend to learn a function that refines already learned features at the same feature level.
 If a residual block mostly performs feature refinement while not changing the level of input features, the performance of the residual network is not significantly affected by dropping the block at test time [40, 41].
-However, in typical residual networks, most residual blocks tend to refine features while learning new level features as well, and, hence, random dropping of residual blocks at test time degrades the performance significantly.
-Therefore, we hypothesize that if some designated residual blocks can be encouraged during training to focus more on feature refinement, then these blocks can be skipped to save computation at marginal loss of prediction accuracy at test time.
-To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two consecutive sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2.
-While $\mathbf{F}^{s}_{base}$ learns feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}$.
+However, in typical residual networks, most residual blocks tend to refine features while learning new level features as well, and, hence, random dropping of residual blocks at test time degrades inference performance significantly.
+Therefore, we hypothesize that if some selected residual blocks can be encouraged explicitly during training to focus more on feature refinement, then these blocks can be skipped to save computation at marginal loss of prediction accuracy at test time.
+To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2.
+We train these two sub-paths to have different properties.
+While $\mathbf{F}^{s}_{base}$ is trained to learn feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}(=\mathbf{h}^{s}_{L+1}$).
 Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference.
-However, since layers in $\mathbf{F}^{s}_{skippable}$ only refine $\mathbf{h}^{s}_{base}$, they can be skipped to save computation.
+In contrast, layers in $\mathbf{F}^{s}_{skippable}$ can be skipped to save computation since they only refine $\mathbf{h}^{s}_{base}$.
 If $\mathbf{F}^{s}_{skippable}$ is skipped, then intermediate features $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage.
-Therefore, the overall network depth can be adjusted by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{s})$ or not.
-In Section 4.1, we show that this architectural pattern for adaptive depth networks is generally applicable to a wide range of residual networks.
-3.2 Skip-Aware Self-Distillation
+Therefore, $2^{N_{r}}$ sub-networks with different accuracy-efficiency trade-offs can be selected from a single network by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{r})$ or not.
+Table 4 in Appendix A.1 provides details on how this pattern is applied to both CNNs and transformers.
+3.2 Training Sub-Paths with Self-Distillation
 Preserving the feature level of $\mathbf{h}^{s}_{base}$ in $\mathbf{F}^{s}_{skippable}$ implies, more specifically, that two feature representations $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$ have similar distributions over training input $\mathbf{X}$.
-If $\mathbf{h}^{s}_{super}$ and $\mathbf{h}^{s}_{base}$ have similar distributions, skipping the sub-path $\mathbf{F}^{s}_{skippable}$ during inference results in minimal internal covariate shifts [25] to the following network layers.
-Kullback-Leibler (KL) divergence measures how different two distributions are over the same random variable, and, hence, we use Equation 2 to measure the similarity of two distributions, $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$, over input $\mathbf{X}$:
-$\underset{x\in\mathbf{X}}{D_{KL}}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ (2)
-Algorithm 1 shows our training method, called skip-aware self-distillation, in which Equation 2 is included in the loss function while the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are jointly trained  [^1] Some frameworks, such as Pytorch’s distributed data parallel (DDP), do not support two consecutive forwards passes followed by a single backward pass: $super.forward\rightarrow base.fowrard\rightarrow loss.backward$.
-For such frameworks, we adapt the algorithm to $super.forward\rightarrow loss_{super}.backward\rightarrow base.forward\rightarrow loss_{sub\_path}.backward$.
-The latter method has similar results and is more memory-efficient.
-.
-In Algorithm 1, the forward function of the proposed adaptive depth networks $\mathbf{M}$ accepts an extra argument, ’skip’, that controls which residual stages skip their skippable sub-paths.
-For example, if $\mathbf{M}$ has 4 residual stages, its base-net is selected by passing ‘skip=[True, True, True, True]’.
+Algorithm 1 shows our training method, in which $\mathbf{h}^{s}_{base}$ and $\mathbf{h}^{s}_{super}$ are encouraged to have similar distributions by including Kullback-Leibler (KL) divergence between them, or ${D_{KL}}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$, in the loss function.
 Algorithm 1 Training algorithm for an adaptive depth network $\mathbf{M}$.
-Two parameter-sharing sub-networks, or super-net and base-net, are jointly trained using skip-aware self-distillation.
-$N_{r}$ denotes the number of residual stages.
+The forward function of $\mathbf{M}$ accepts an extra argument, ‘skip’, which controls the residual stages where their skippable sub-paths are skipped.
+For example, the smallest sub-network, or base-net, of $\mathbf{M}$ is selected by passing ‘skip=[True, True, True, True]’ when the total number of residual stages, denoted by $N_{r}$, is 4.
 1:Initialize an adaptive depth network $\mathbf{M}$
@@
 6:     $loss_{super}=criterion(\mathbf{y},\mathbf{\hat{y}}_{super})$
+7:     $loss_{super}.backward()$
 […]
-8:     $loss_{base}=\sum_{s=1}^{N_{r}}D_{KL}(\mathbf{h}^{s}_{super}\|\mathbf{h}^{s}_{base})\newline \hskip 50.00008pt+D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$
-9:     $loss=\alpha loss_{super}+(1-\alpha)loss_{base}$
-10:     $loss.backward()$
+9:     $loss_{base}=\sum_{s=1}^{N_{r}}D_{KL}(\mathbf{h}^{s}_{super}\|\mathbf{h}^{s}_{base})+D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$
+10:     $loss_{base}.backward()$triangleright self-distillation of skippable sub-paths
 11:     $optimizer.step()$
 12:end for
-In steps 5 and 7, the forward passes of the super-net and the base-net are executed for the same input $\mathbf{x}$.
-Intermediate features $\mathbf{h_{super}}$ and $\mathbf{h_{base}}$ are also obtained during the forward passes of $\mathbf{M}$.
-In step 8, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in Equation 2 is included in the loss function $loss_{base}$.
-By minimizing $loss_{base}$, two feature representations $\mathbf{h}^{s}_{super}$ and $\mathbf{h}^{s}_{base}$, are explicitly enforced to have similar distributions for the same input $\mathbf{x}$.
-Further, this step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage.
+In Algorithm 1, the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are exploited.
+During the forward passes of the super-net and the base-net in steps 5 and 8, respectively, intermediate features $\mathbf{h}_{super}$ and $\mathbf{h}_{base}$ are obtained.
+In steps 9-10, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in the loss function is explicitly minimized for training data $\mathbf{X}$.
+Further, this self-distillation step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage.
 Therefore, $\mathbf{h}^{s}_{base}$ is expected to learn more compact representation from $\mathbf{h}^{s}_{super}$.
-The hyperparameter $\alpha$ controls the strength of this skip-aware self-distillation.
-In step 8, $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is included in the loss function for further distillation effect from the super-net to the sub-networks.
-Due to the architectural pattern of interleaving the mandatory and the skippable sub-paths, minimizing $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ also minimizes $D_{KL}(\mathbf{h}_{super}\|\mathbf{h}_{base})$ implicitly.
-Experimental result shows that similar results can be achieved when only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is used in $loss_{base}$ (Section 4.5).
-This implicit approach is useful when the extraction of intermediate features is tricky.
-In Algorithm 1, only two sub-networks are involved for the training of $\mathbf{M}$, and, hence, the total training time is no greater than training two sub-networks individually.
-However, at test time, sub-networks with various depths can be selected on the fly by systematically skipping sub-paths in residual stages.
-For example, in a network with 4 residual stages, $16\ (=2^{4})$ parameter-sharing sub-networks can be selected by varying the skip argument.
-In contrast, prior adaptive networks supporting 16 parameter-sharing sub-networks need to perform 16 explicit self-distillation from the super-net to the sub-networks [48, 59, 20].
-We demonstrate this result in Section 4.3.
-3.3 Formal Analysis of Skippable Sub-Paths
-$D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in Equation 2 can be trivially minimized if residual blocks in $\mathbf{F}^{s}_{skippable}$ learn identity functions, or $\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})=\mathbf{h}^{s}_{base}$.
+Self-distillation itself has been extensively used in prior adaptive networks [11, 10, 13].
+However, their primary goal is to train every target sub-network.
+For example, on each iteration, sub-networks are randomly sampled from a large search space to act as either teachers or students [11], which results in significantly longer training time.
+In contrast, our method in Algorithm 1 focuses on training sub-paths by exclusively using the super-net as a teacher and the base-net as a student.
+By focusing on training sub-paths rather than target sub-networks, the training procedure in Algorithm 1 is significantly simplified.
+At test time, however, $2^{N_{r}}(=\sum_{k=0}^{N_{r}}{C(N_{r},k)})$ different sub-networks can be constructed instantly by connecting these sub-paths in a combinatorial manner.
+The effect of this self-distillation strategy is investigated in Section 4.4.
+3.3 Analysis of Skippable Sub-Paths
+Formal Analysis: $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in the loss function $loss_{base}$ can be trivially minimized if residual blocks in $\mathbf{F}^{s}_{skippable}$ learn identity functions, or $\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})=\mathbf{h}^{s}_{base}$.
 However, since the super-net is jointly trained with the loss function $loss_{super}$, the residual functions in $\mathbf{F}^{s}_{skippable}$ cannot simply be an identity function.
@@
 $\mathcal{L}(\mathbf{h}^{s}_{super})=\mathcal{L}\{\mathbf{h}^{s}_{base}+\mathbf{F}^{s}_{skippable}(\mathbf{h}^{s}_{base})\}$ (2)
-=$=\mathcal{L}\{\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})+F_{L}(\mathbf{h}^{s}_{L})\}$ (4)
-$\approx\mathcal{L}\{\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})\}$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +F_{L}(\mathbf{h}^{s}_{L})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{L})}{\partial\mathbf{h}^{s}_{L}}+\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ (5)
+=$=\mathcal{L}\{\mathbf{h}^{s}_{base}+F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})+...+F_{L-1}(\mathbf{h}^{s}_{L-1})+F_{L}(\mathbf{h}^{s}_{L})\}$ (3)
+$\approx\mathcal{L}\{\mathbf{h}^{s}_{base}+F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})+...+F_{L-1}(\mathbf{h}^{s}_{L-1})\}+F_{L}(\mathbf{h}^{s}_{L})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{L})}{\partial\mathbf{h}^{s}_{L}}+\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ (4)
 In Equation 4, the loss function is expanded around $\mathbf{h}^{s}_{L}$, or $\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})$.
 Only the first order term is left and all high order terms, such as $F_{L}(\mathbf{h}_{L}^{s})^{2}\cdot\frac{{\partial^{2}\mathcal{L}(\mathbf{h}^{s}_{L})}}{2{\partial(\mathbf{h}^{s}_{L})^{2}}}$, are absorbed in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$.
-The terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude.
+The high-order terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude.
 In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude.
-In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the skip-aware self-distillation, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation.
-If we similarly keep expanding the loss function around $\mathbf{h}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
+In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the proposed self-distillation strategy, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation; see Figure 3 for empirical evidence.
+If we similarly keep expanding the loss function around $\mathbf{h}^{s}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
 $\mathcal{L}(\mathbf{h}^{s}_{super})\approx\mathcal{L}(\mathbf{h}^{s}_{base})+\sum_{j={L/2+1}}^{L}F_{j}(\mathbf{h}^{s}_{j})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{j})}{\partial\mathbf{h}^{s}_{j}}$ (5)
@@
 In other words, the residual functions in the skippable sub-paths reduce the loss $\mathcal{L}(\mathbf{h}^{s}_{base})$ iteratively during inference while preserving the feature distribution of $\mathbf{h}^{s}_{base}$.
-Considering this result, we can conjecture that, with our architectural pattern and training method, layers in $\mathbf{F}^{s}_{skippable}$ learn functions that refine input features $\mathbf{h}^{s}_{base}$ iteratively for better inference accuracy while minimally changing the distribution of $\mathbf{h}^{s}_{base}$.
-Therefore, skipping $\mathbf{F}^{s}_{skippable}$ only slightly reduces prediction accuracy.
-3.4 Skip-aware Batch Normalization
+Considering this result, we can conjecture that, with our architectural pattern and self-distillation strategy, layers in $\mathbf{F}^{s}_{skippable}$ learn functions that refine input features $\mathbf{h}^{s}_{base}$ iteratively for better inference accuracy while minimally changing the distribution of $\mathbf{h}^{s}_{base}$.
+Figure 3: $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at residual blocks. In ours, skippable sub-paths (orange areas) minimally change the distribution of input h
+Empirical Analysis: We can estimate how much the distribution of input $\mathbf{h}$ is transformed by residual function $F$ by measuring $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at each residual block.
+Figure 3 illustrates $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at every residual block in the baseline ResNet50 and our ResNet50-ADN that is trained according to Algorithm 1  [^2] ImageNet validation dataset is used for this experiment..
+Our ResNet50-ADN exhibits greater transformation than ResNet50 in the mandatory sub-paths (blue areas) and less transformation in the skippable sub-paths (orange areas).
+This result demonstrates that our self-distillation strategy in Algorithm 1 effectively trains the skippable sub-paths to minimally change the input distribution.
+As a result, the blocks in skippable sub-paths can be skipped with less impact on performance.
+3.4 Skip-Aware Batch/Layer Normalization
 Originally, batch normalization (BN) [45] was proposed to handle internal covariate shift during training non-adaptive networks by normalizing features.
@@
 4 Experiments
-To demonstrate the generality and effectiveness of our approach for adaptive depth networks, we conduct experiments on various networks and vision tasks.
-4.1 Networks
-We use four representative residual networks as base models to apply the proposed architecture pattern in Section 3.1: MobileNet V2 [38] is a lightweight CNN model, ResNet [16] is a larger CNN model, and ViT [5] and Swin-T [34] are representative vision transformers.
-All base models except ViT have 4 residual stages, each with 2 $\sim$ 6 (residual or encoder) blocks.
+We use four representative residual networks both from CNNs and vision transformers as base models to apply the proposed architecture pattern: MobileNet V2 [35] is a lightweight CNN model, ResNet [34] is a larger CNN model, and ViT [37] and Swin-T [38] are representative vision transformers.
+All base models except ViT have hierarchical stages, each with 2 $\sim$ 6 residual blocks.
 So, according to the proposed architectural pattern, every residual stage is evenly divided into 2 sub-paths for depth adaptation.
-Since ViT does not define residual stages, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and desiginate the last encoder block of each group as a skippable sub-path.
+ViT does not define hierarchical stages and all residual encoder blocks have same spatial dimensions.
+Therefore, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and select the last encoder block of each group as a skippable sub-path.
+Details are in Table 4 in Appendix A.1.
 We use the suffix ‘-ADN’ to denote our adaptive depth networks.
@@
 For example, ResNet50-ADN (base-net) is equivalent to ResNet50-ADN (TTTT).
-Table 1: Results on ImageNet.
-Individual networks with the suffix ‘-Base’ have the same depths as the base-nets of corresponding adaptive depth networks.
-Latency is measured on an RTX 3090 (batch size = 64).
-Model Params FLOPs Acc@1 Latency
-(M) (G) (%) (msec)
-ResNet50-ADN (super-net) 25.58 4.11 77.6 43.2
-ResNet50-ADN (base-net) 2.58 76.1 27.6
-ResNet50 (individual) 25.56 4.11 76.7 43.2
-ResNet50-Base (individual) 17.11 2.58 75.0 27.6
-MV2-ADN (super-net) 3.72 0.32 72.7 15.5
-MV2-ADN (base-net) 0.25 70.7 13.2
-MV2 (individual) 3.50 0.32 72.1 15.5
-MV2-Base (individual) 2.99 0.25 70.7 13.2
-ViT-b/16-ADN (super-net) 86.59 17.58 81.3 138.6
-ViT-b/16-ADN (base-net) 11.76 79.4 90.7
-ViT-b/16 (individual) 86.57 17.58 81.1 138.6
-ViT-b/16-Base (individual) 67.70 11.76 78.7 90.7
-Swin-T-ADN (super-net) 28.30 4.49 81.8 74.4
-Swin-T-ADN (base-net) 2.69 78.7 39.9
-Swin-T (individual) 28.29 4.49 81.5 74.4
-Swin-T-Base (individual) 15.63 2.69 78.6 39.9
+Model Params FLOPs Acc
+(M) (G) (%)
+ResNet50-ADN (FFFF) 25.58 4.11 77.6
+ResNet50-ADN (TTTT) 2.58 76.1
+ResNet50 25.56 4.11 76.7
+ResNet50-Base 17.11 2.58 75.0
+MbV2-ADN (FFFF) 3.72 0.32 72.7
+MbV2-ADN (TTTT) 0.25 70.7
+MbV2 3.50 0.32 72.1
+MbV2-Base 2.99 0.25 70.7
+ViT-b/16-ADN (FFFF) 86.59 17.58 81.3
+ViT-b/16-ADN (TTTT) 11.76 79.4
+ViT-b/16 86.57 17.58 81.1
+ViT-b/16-Base 67.70 11.76 78.7
+Swin-T-ADN (FFFF) 28.30 4.49 81.8
+Swin-T-ADN (TTTT) 2.69 78.7
+Swin-T 28.29 4.49 81.5
+Swin-T-Base 15.63 2.69 78.6
+Figure 4: (a) Results on ImageNet validation dataset.
+Networks with the suffix ‘-Base’ have the same depths as the base-nets of corresponding adaptive depth networks.
+(b) Pareto frontiers formed by the sub-networks of our adaptive depth networks.
+ResNet50 (individual) and ResNet50 (KD individual) are non-adaptive networks having same depths as the sub-networks of ResNet50-ADN.
 4.1 ImageNet Classification
@@
 ViT and Swin-T are trained for 300 epochs, following DeiT’s training recipe [50, 51].
-However, in Swin-T-ADN, we disable stochastic depths [23] for the base-net since the strategy of random dropping of residual blocks conflicts with our approach to skipping sub-paths.
+For Swin-T-ADN, we disable stochastic depths [40] for the mandatory sub-paths since the strategy of random dropping of residual blocks conflicts with our approach to skipping sub-paths.
 For fair comparison, our adaptive depth networks and corresponding individual networks are trained in the same training settings.
-The hyperparameter $\alpha$ in Algorithm 1 is set to $0.5$ for all networks.
-The results in Table 1 show that our adaptive depth networks outperform individual counterpart networks even though our sub-networks share parameters in a single network.
-Further, our results with vision transformers demonstrate that our approach is generally applicable to residual networks and compatible with their state-of-the-art training techniques such as DeiT’s training recipe [42].
+The results in Figure 4-(a) show that our adaptive depth networks outperform counterpart individual networks even though many sub-networks share parameters in a single model.
+Further, our results with vision transformers demonstrate that our approach is generally applicable and compatible with their state-of-the-art training techniques such as DeiT’s training recipe [50].
 We conjecture that this performance improvement results from effective distillation of knowledge from $\textbf{h}^{s}_{super}$ to $\textbf{h}^{s}_{base}$ at each residual stage and the iterative feature refinement at skippable sub-paths, shown in Equation 6.
-In Table 2 and Figure 3, several state-of-the-art efficient inference methods and dynamic networks are compared with our sub-networks of adaptive depth networks.
-The result demonstrates that our adaptive depth networks match or outperform many state-of-the-art static and dynamic networks across varying depth ranges.
-In Figure 3, sub-networks of ResNet50-ADN are compared to the counterpart individual ResNets of equivalent depths.
-In particular, it should be noted that ResNets trained with knowledge distillation in the same training settings has worse performance than individual ResNets trained without knowledge distillation.
+Figure 4-(b) shows Pareto frontiers formed by selected sub-networks of our adaptive depth networks; Table 5 in Appendix A.2 shows the performance of all sub-networks.
+In Figure 4-(b) and Table 1, several state-of-the-art efficient inference methods and dynamic networks are compared with our base-networks.
+The result demonstrates that our adaptive depth networks show competitive performance across a range of FLOPs.
+In Figure 4-(b), it should be noted that individual ResNets trained with knowledge distillation has worse performance than individual ResNets.
 As reported in previous works, successful knowledge distillation requires a patient and long training [52], and straightforward knowledge distillation using ImageNet does not improve the performance of student models [53, 54].
-In contrast, our ResNet50-ADN trained with skip-aware self-distillation achieves better performance than counterpart ResNets.
-This demonstrates that the high performance of adaptive depth networks does not simply come from the distillation effect, but from the effective combination of the proposed architectural pattern and the skip-aware self-distillation strategy.
-Figure 3: Comparison with various approaches on ImageNet.
-Sub-networks of our adaptive depth networks match or outperform other approaches on ImageNet.
-ResNet50 (individual) shows individual networks that have equivalent network depths as the sub-networks of ResNet50-ADN.
-ResNet50 (KD individual) shows individual networks trained using knowledge distillation from pretrained Pytorch ResNet50.
-In legend, $\dagger$ denotes static compression methods, and $\star$ denotes input-dependent dynamic networks.
-Table 2: Our base-nets are compared with state-of-the-art efficient inference methods for various networks.
-$\dagger$ denotes static compression methods, $\ast$ denotes width-adaptation networks, and $\star$ denotes input-dependent dynamic networks.
-Note that many of these state-of-the-art methods exploit various non-canonical training techniques, such as iterative retraining and post-processing, while our base-nets are selected from adaptive depth networks without any fine-tuning.
+In contrast, our ResNet50-ADN trained with the proposed self-distillation strategy achieves better performance than counterpart ResNets.
+This demonstrates that the high performance of adaptive depth networks does not simply come from distillation effect.
+Table 1: Our base-nets are compared with state-of-the-art efficient inference methods.
+$\dagger$ denotes static pruning methods, $\ast$ denotes width-adaptation networks, and $\star$ denotes input-dependent dynamic networks.
+While these approaches exploit various non-canonical training techniques, such as iterative retraining, our base-nets are instantly selected from adaptive depth networks without fine-tuning.
 […]
-GReg-1 [49$\dagger$ 2.8G 33% 76.1%
-DR-ResNet50 ($\alpha$=2.0) [62$\star$ 2.3G 44% 75.3%
-DepGraph [7$\dagger$ 2.0G 51% 75.8%
-ResNet50-ADN (base-net) 2.6G 37% 76.1%
+GReg-1 [55$\dagger$ 2.8G 33% 76.1%
+DepGraph [56$\dagger$ 2.0G 51% 75.8%
+DR-ResNet50 ($\alpha$=2.0) [47$\star$ 2.3G 44% 75.3%
+ResNet50-ADN (TTTT) 2.6G 37% 76.1%
 AlphaNet-0.75x [11] $\ast$ 0.21G 34% 70.5%
-MV2-ADN (base-net) 0.25G 22% 70.7%
-Ada-DeiT-S [35$\star$ 2.3G - 77.3%
-WDPruning-0.3-11 [56$\dagger$ 2.6G - 78.3%
-DeiT-S+ATS (not fine-tuned) [8$\star$ 3.0G - 78.5%
-SPViT-Swin-Ti [28$\star$ 3.3G 27% 80.1%
-Swin-T-ADN (base-net) 2.7G 40% 78.7%
+MbV2-ADN (TTTT) 0.25G 22% 70.7%
+WDPruning-0.3-11 [57$\dagger$ 2.6G - 78.3%
+X-Pruner [58] $\dagger$ 3.2G 28% 80.7%
+Ada-DeiT-S [27$\star$ 2.3G - 77.3%
+SPViT-Swin-Ti [59$\star$ 3.3G 27% 80.1%
+Swin-T-ADN (TTTT) 2.7G 40% 78.7%
 4.2 Training Time
-Table 3: Training time (1 epoch), measured on a Nvidia RTX 4090 (batch size = 128). All compared models are based on ResNet50 or have similar FLOPs and depths.
-Model Training time (min) # of sub-networks
-ResNet50 32.9 1
-ResNet50-Base 22.7 1
-ResNet50-ADN (ours) 54.7 $2^{4}$ (varying depths)
-MSDNet [24] 67.1 9 (early-exit branches)
-S-ResNet50 [59] 82.5 4 (varying widths)
-SkipNet [50] 90.6 dynamic depths
-One important advantage of our approach is that our adaptive depth networks require significantly shorter training time than other adaptive networks.
-Table 3 shows that training our ResNet50-ADN takes similar time to train two individual networks combined.
-In contrast, the compared adaptive networks requires much longer training time than ours.
-MSDNet [24] represents adaptive depth networks with multiple early-exiting branches and classifiers.
-SkipNet [50] represents input-dependent dynamic networks that drop layers for ‘easy’ inputs.
-In S-ResNet50 [59], 4 sub-networks with varying widths are trained through iterative self-distillation from its super-net.
-Despite significantly shorter training time is required, our adaptive depth networks can support various sub-networks at test time.
-Figure 4-(left) shows the performance of ResNet50-ADN when its depth is varied at test time.
-Among them, only the super-net and the base-net are explicitly trained in Algorithm 1.
-Other sub-networks are selected on the fly at test time by skipping sub-paths in varying number of residual stages.
-Although these sub-networks are not trained explicitly, they show graceful degradation of performance as the depth of sub-networks becomes gradually shallower.
-However, although ResNet50-ADN support $2^{4}$ sub-networks in theory, Figure 4-(right) shows that some sub-networks are more useful than the others.
-For instance, even though ResNet50-ADN(TFFF) and ResNet50-ADN(FFFT) have the same network depths, ResNet50-ADN(TFFF) has about $0.58\%$ higher top-1 accuracy.
-This result shows that skipping sub-paths in later stages is more detrimental to the performance in general.
-Figure 4: Validation accuracy of our sub-networks of varying depths. The boolean values below the bars indicate the residual stages where the sub-paths are skipped.
-4.4 Performance on Devices
-Figure 5 shows the performance of adaptive networks in an actual device.
-The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50 [59], a representative width-adaptation network.
+One major advantage of our adaptive depth networks is that they require significantly shorter training time than prior adaptive networks.
+Figure 5-(a) shows that training our adaptive depth networks take similar time to train two individual networks combined.
+This is because our training method in Algorithm 1 exclusively exploits only the super-net and the base-net for self-distillation.
+In contrast, compared adaptive networks require much longer training time than ours since they have to train every target sub-network explicitly  [^3] Although the prior works we compared for training time may seem outdated, they remain relevant and representative because there has been little progress in reducing the training time of adaptive networks..
+For example, on every training iteration, AlphaNet [11] randomly samples sub-networks from its search space for self-distillation.
+While AlphaNet [11] supports significantly more sub-networks, it should be noted that supporting many sub-networks is not our goal.
+Instead, our objective is to provide better performance Pareto with several useful sub-networks, as demonstrated in Figure 4-(b).
+For example, although ResNet50-ADN can select $2^{4}$ sub-networks, Table 5 (in Appendix A.2) shows that some sub-networks are more useful than the others; results in Table 5 show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
+Model 1 epoch (min) # of sub-nets
+ResNet50 32.9 -
+ResNet50-Base 22.7 -
+ResNet50-ADN (ours) 54.7 $2^{4}$ depths
+MSDNet [8] 67.1 9 depths
+S-ResNet50 [10] 82.5 4 widths
+MbV2 13.1 -
+MbV2-Base 10.0 -
+MbV2-ADN (ours) 22.5 $2^{4}$ depths
+AlphaNet [11] 230.5 216 depths
+Figure 5: (a) Training time (1 epoch), measured on Nvidia RTX 4090 (batch size: 128).
+AlphaNet is configured to have similar FLOPs to MbV2 and makes sub-networks by only adjusting the network depth.
+(b) Inference latency and energy consumption of adaptive networks, measured on Nvidia Jetson Orin Nano (batch size: 1)
+4.3 On-Device Performance
+While the Pareto frontiers formed by sub-networks demonstrate theoretic performance, inference acceleration in actual devices is more important in practice for effective control of inference latency and energy consumption.
+Figure 5(b) shows the performance in Nvidia Jetson Orin Nano.
+The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50, a representative width-adaptation network.
 The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption.
@@
 In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
-Figure 5: Inference latency and energy consumption of adaptive networks, measured on Nvidia Jetson Orin Nano (batch size = 1).
-Although adjusting the network width can support a much wider FLOPs adaptation range, depth adaptation of ResNet50-ADN is more effective for actual inference acceleration and saving energy.
-Table 4: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN.
-Applied components are checked.
-In the last row, only $D_{KL}$ with softmax outputs is used for the skip-aware self-distillation.
-$\downarrow$ and $\uparrow$ in parentheses are comparisons to individual networks.
-Skip-aware Skip-aware ResNet50 Acc@1 ViT-b/32 Acc@1
-self-distillation BNs / LNs Super-net Base-net Super-net Base-net
+4.4 Ablation Study
+Table 2: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. $\downarrow$ and $\uparrow$ in parentheses are comparisons to non-adaptive individual networks.
+self- skip-aware ResNet50 Acc@1 (%) ViT-b/32 Acc@1 (%)
+distllation. BNs/LNs FFFF TTTT FFFF TTTT
 75.2% ($\downarrow$ 1.5%) 72.2% ($\downarrow$ 2.8%) 75.7% ($\downarrow$0.2%) 74.1% ($\uparrow$ 0.3%)
@@
 O O 77.6% ($\uparrow$ 0.9%) 76.1% ($\uparrow$ 1.1%) 76.6% ($\uparrow$ 0.8%) 74.3% ($\uparrow$0.5%)
-$\triangle$ O 77.6% ($\uparrow$ 0.9%) 75.9% ($\uparrow$ 0.9%)
-4.5 Ablation Study
-We conduct ablation study on ImageNet classification to investigate the influence of two key components of the proposed adaptive depth networks: (1) skip-aware self-distillation and (2) skip-aware BN/LNs.
-When our skip-aware self-distillation is not applied, the loss in Algorithm 1 is modified to $loss=\frac{1}{2}\{criterion(\mathbf{y},\mathbf{\hat{y}}_{super})+criterion(\mathbf{y},\mathbf{\hat{y}}_{base})\}$ for joint training of the super-net and the base-net.
+We first investigate the influence of two key components of the proposed adaptive depth networks: (1) self-distillation of sub-paths and (2) skip-aware BNs/LNs.
+When our self-distillation method is not applied, the loss of the base-net, or $loss_{base}$, in Algorithm 1 is modified to $criterion(\mathbf{y},\mathbf{\hat{y}}_{base})$.
 Table 2 shows the results.
-For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than individual networks by 1.5% and 2.8%, respectively.
+For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than non-adaptive individual networks by 1.5% and 2.8%, respectively.
 This result shows the difficulty of joint training sub-networks for adaptive networks.
 When one of the two components is applied individually, the performance is still slightly worse than individual networks’.
-In the third row, when both skip-aware self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net.
-Finally, the last row shows the result when only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is used for skip-aware self-distillation.
-Without explicit distillation of intermediate features, slightly lower performance is observed in the base-net.
+In the last row, when both self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net.
 The results with ViT-b/32 show that switchable layer normalization has similar effect in vision transformers.
-(a) Original ResNet50
-(b) ResNet50-ADN (super-net)
-(c) ResNet50-ADN (base-net)
-Figure 6: Class Activation Maps of the 3rd residual stages of ResNet50s.
-(a) Original ResNet50’s activation regions change gradually across all blocks.
-(b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features.
-In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target.
-(c) The base-net’s final activation map is very similar to super-net’s final activation map, implying that they are at the same feature level.
-4.6 Visual Analysis of Sub-Paths
-To investigate how our training method affects feature representations in the mandatory and the skippable sub-paths, we visualize the activation of 3rd residual stage of ResNet50-ADN using Grad-CAM [39].
-The 3rd residual stage of ResNet50-ADN has 6 residual blocks and the last three blocks are skippable.
-In Figure 6-(a), the activation regions of original ResNet50 changes gradually across all consecutive blocks.
-In contrast, in Figure 6-(b), ResNet50-ADN (super-net) manifests very different activation regions in two sub-paths.
-In the first three residual blocks, we can observe lots of hot activation regions in wide areas, suggesting active learning of new level features.
-In contrast, significantly less activation regions are found in the skippable last three blocks and they are gradually concentrated around the target object, demonstrating the refinement of learned features.
-Further, in Figure 6-(c), we can observe that the final activation map of the ResNet50 (base-net) is very similar to the super-net’s final activation map in Figure 6-(b).
-This implies that they have similar distributions for the same inputs, as suggested in Section 3.3.
-4.7 Object Detection and Instance Segmentation
+Table 3: Comparison of self-distillation strategies. Our approach (in bold) uses exclusively the super-net and the base-net as a teacher and a student, respectively.
+sub-nets ResNet50-ADN Acc@1 (%) ViT/b-32-ADN Acc@1 (%)
+Teacher Student FFFF TFFF TTFF TTTF TTTT FFFF TFFF TTFF TTTF TTTT
+FFFF TTTT 77.6 77.2 76.8 76.1 76.1 76.6 76.0 75.5 74.5 74.3
+FFFF Random 77.1 76.7 76.4 75.5 74.8 75.2 74.7 73.9 72.9 71.1
+Random TTTT 75.5 75.5 75.4 75.0 74.9 72.0 72.0 71.9 71.8 71.7
+Random Random 75.4 75.2 75.2 74.9 74.6 70.8 70.7 70.6 70.4 70.3
+Self-Distillation Strategies: Our self-distillation approach in Algorithm 1 exploits only two sub-networks exclusively as a teacher and a student.
+Specifically, the super-nets, or FFFF, acts as the the teacher and the base-nets, or TTTT, becomes the student.
+The purpose of this strategy is not to train only those two sub-networks, but rather to train skippable sub-paths in a way that minimally modifies the feature distribution, as demonstrated in Figure 3.
+To investigate the effect of self-distillation strategies, we conduct an experiment in Table 3.
+In every training iteration, rather than exclusively using FFFF and TTTT sub-networks for self-distillation, we randomly sample either a teacher, a student, or both from $2^{4}$ sub-networks.
+These randomly sampled sub-networks are trained explicitly through self-distillation.
+However, as shown in Table 3, all sub-networks, both from ResNet50-ADN and ViT/b-32-ADN, perform significantly better when our self-distillation strategy is applied for training.
+Even though some sub-networks (such as TTTF, TTFF, and TTTF) are not trained explicitly with our self-distillation strategy, they still outperform their counterpart sub-networks trained through random sampling.
+This result demonstrates that our method of training sub-paths is more effective than training target sub-networks themselves.
+Figure 6: Pareto frontiers of ResNet50-ADNs trained with varying ratios between mandatory and skippable sub-paths. Total number of blocks remains unchanged.
+Lengths of Mandatory Sub-Paths: In Figure 6, we investigate the effect of varying the length ratio between the mandatory and the skippable sub-paths.
+(Details are in Table 6.) If mandatory sub-paths become shorter than skippable sub-paths, smaller sub-networks can be selected since more layers can be skipped.
+However, the result in Figure 6 shows that this configuration significantly degrades performance of all sub-networks.
+Since every sub-network shares parameters of mandatory sub-paths, low inference capability of shallow mandatory sub-paths affects all sub-networks.
+This implies that maintaining certain depths in mandatory sub-paths is essential for effective inference.
+In contrast, increasing the length of mandatory sub-paths further does not yield further performance improvement and only reduce the range of FLOPs adaptation.
+5 Conclusions
+We propose a practical approach to adaptive depth networks that can be applied to both CNNs and transformers with minimal training effort.
+We provide a general principle and formal explanation on how depth adaptation can be achieved with minimal performance degradation.
+Under this principle, our approach can avoid typical exhaustive training of target sub-network and instead focus on optimizing the sub-paths of the network to have specific properties.
+At test time, those sub-paths can be connected in a combinatorial manner to construct sub-networks with various accuracy-efficiency trade-offs from a single network.
+Experimental results show that those sub-networks form a better Pareto frontier than non-adaptive baseline networks and achieve actual inference acceleration.
+Appendix A Appendix: Detailed Settings and Evaluation Results
+A.1 Detailed Architectures
+number of number of
+mandatory skippable total
+blocks blocks blocks
+stage 1 2 1 3
+stage 2 2 2 4
+stage 3 3 3 6
+stage 4 2 1 3
+(a) ResNet50-ADN
+number of number of
+mandatory skippable total
+blocks blocks blocks
+stage 1 2 1 3
+stage 2 2 1 3
+stage 3 2 1 3
+stage 4 2 1 3
+(b) Vit-b/16-ADN
+number of number of
+mandatory skippable total
+blocks blocks blocks
+stage 1 1 - 1
+stage 2 2 - 2
+stage 3 2 1 3
+stage 4 2 2 4
+stage 5 2 1 3
+stage 6 2 1 3
+stage 7 1 - 1
+(c) MbV2-ADN
+number of number of
+mandatory skippable total
+blocks blocks blocks
+stage 1 1 1 2
+stage 2 1 1 2
+stage 3 3 3 6
+stage 4 2 - 2
+(d) Swin-T-ADN
+Table 4: Each residual stage of base models is evenly divided into two sub-paths; the first is mandatory and the other is skippable.
+Since ViT does not define hierarchical stages, 12 identical encoder blocks are divided into 4 stages, resembling other residual networks for vision tasks.
+A.2 Performance of Sub-Networks
+sub-network FLOPs (G) Acc@1 (%)
+FFFF 4.11 77.6
+TFFF 3.90 77.2
+FTFF 3.68 76.5
+FFTF 3.46 75.6
+FFFT 3.90 76.2
+TTFF 3.46 76.8
+TFTF 3.24 75.4
+TFFT 3.68 76.1
+FTTF 3.02 75.3
+FTFT 3.46 75.4
+FFTT 3.25 75.6
+TTTF 2.80 76.1
+TTFT 3.24 75.9
+TFTT 3.02 75.4
+FTTT 2.80 75.6
+TTTT 2.58 76.1
+(a) ResNet50-ADN
+sub-network FLOPs (G) Acc@1 (%)
+FFFF 17.58 81.3
+TFFF 16.20 80.5
+FTFF 16.20 80.2
+FFTF 16.20 80.0
+FFFT 16.20 80.4
+TTFF 14.67 79.9
+TFTF 14.67 79.6
+TFFT 14.67 79.8
+FTTF 14.67 79.4
+FTFT 14.67 79.7
+FFTT 14.67 79.4
+TTTF 13.21 79.3
+TTFT 13.21 79.4
+TFTT 13.21 79.2
+FTTT 13.21 79.1
+TTTT 11.76 79.1
+(b) ViT-b/16-ADN
+Table 5: FLOPs and ImageNet validation accuracy of sub-networks.
+Only the super-net and the base-net are trained explicitly.
+The other parameter-sharing sub-networks in the middle can be selected at test time without explicit training.
+The highest accuracy in each group is shown in bold.
+This results show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
+A.3 Varying the Lengths of Sub-Paths in ResNet50-ADN
+number of number of total
+mandatory skippable blocks
+blocks blocks
+stage 1 1 2 3
+stage 2 1 3 4
+stage 3 2 4 6
+stage 4 1 2 3
+(a) # of mandatory $<$ # of skippable
+number of number of total
+mandatory skippable blocks
+blocks blocks
+stage 1 2 1 3
+stage 2 3 1 4
+stage 3 4 2 6
+stage 4 2 1 3
+(b) # of mandatory $>$ # of skippable
+Table 6: Varying the length ratio of sub-paths of ResNet50-ADN. Total number of blocks of each stage is not changed.
+A.4 Varying Sub-Paths Length Ratios in Vit-b/16-ADN
+Each stage of Vit-b/16-ADN has 3 encoder blocks and, by default, we select every last encoder block as a skippable sub-path.
+Therefore every stage has two mandatory blocks and 1 skipable blocks as shown in Table 4-(b).
+Figure 7-(a) shows different configuration where every last two blocks of the stages become skippable.
+With this configuration in Figure 7-(a), we can select much smaller sub-networks.
+For example, the smallest sub-network, or the base-net, of Vit-b/15-ADN has only 4 mandatory blocks and it requires 5.82 GFLOPs.
+However, the result in Figure 7-(b) shows that this configuration significantly degrades performance of all sub-networks.
+As demonstrated with ResNet50-ADN in Figure 6, low inference capability of shallow mandatory sub-paths affects all sub-networks.
+This again shows that some minimum depths in mandatory sub-paths is required for effective inference.
+number of number of total
+mandatory skippable blocks
+blocks blocks
+stage 1 1 2 3
+stage 2 1 2 3
+stage 3 1 2 3
+stage 4 1 2 3
+Figure 7: (a) The configuration of Vit-b/16-ADN with longer skippable sub-paths. (b) Pareto-frontier when different length ratios between the mandatory and the skippable sub-paths are applied.
+Appendix B Appendix: More Experiments and Analysis
+B.1 Object Detection and Instance Segmentation
 Table 7: Object detection and instance segmentation results on MS COCO dataset.
@@
 Our adaptive depth backbone networks still outperform individual static backbone networks in terms of COCO’s standard metric AP.
-5 Conclusions
-We propose an architectural pattern for adaptive depth networks and its training method that renders selected sub-paths of the network to incur minimal performance degradation even if they are skipped during inference.
-Our approach does not train a fixed set of sub-networks iteratively, and, hence significantly shorter training time is required than previous adaptive networks.
-However, once deployed on devices, it can instantly construct sub-networks of various depths to provide various accuracy-efficiency trade-offs in a single network.
-We show formally that the proposed architectural pattern and training method reduces overall prediction errors while minimizing the impact of skipping sub-paths.
-We also empirically demonstrate the generality and effectiveness of our approach using representative convolutional neural networks and vision transformers.
+B.2 Visual Analysis of Sub-Paths
+To investigate how our training method affects feature representations in the mandatory and the skippable sub-paths, we visualize the activation of 3rd residual stage of ResNet50-ADN using Grad-CAM [63].
+The 3rd residual stage of ResNet50-ADN has 6 residual blocks and the last three blocks are skippable.
+In Figure 8-(a), the activation regions of original ResNet50 changes gradually across all consecutive blocks.
+In contrast, in Figure 8-(b), ResNet50-ADN(FFFF), or super-net, manifests very different activation regions in two sub-paths.
+In the first three residual blocks, we can observe lots of hot activation regions in wide areas, suggesting active learning of new level features.
+In contrast, significantly less activation regions are found in the skippable last three blocks and they are gradually concentrated around the target object, demonstrating the refinement of learned features.
+While ResNet50-ADN(TTTT), or base-net, shares the parameters with the super-net in the first 3 mandatory blocks, their activation regions are very different from the super-net’s.
+This is because while the super-net and the base-net share parameters in the non-skippable mandatory blocks, they use different batch normalization operators in the mandatory sub-paths.
+Further, in Figure 8-(c), we can observe that the final activation map of the base-net is very similar to the super-net’s final activation map in Figure 8-(b).
+This implies that they have similar distributions for the same inputs, as suggested in Section 3.3.
+(a) Original ResNet50
+(b) ResNet50-ADN (FFFF), or super-net
+(c) ResNet50-ADN (TTTT), or base-net
+Figure 8: Class Activation Maps of the 3rd residual stages of ResNet50s.
+(a) Original ResNet50’s activation regions change gradually across all blocks.
+(b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features.
+In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target.
+(c) Even though parameters are shared, the activation map of base-net is very different from super-net’s since they use different batch normalization operators.

### v3

@@
 However, previous adaptive depth networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps.
-In this paper, we present a practical approach to adaptive depth networks that is applicable to various networks with minimal training effort.
+In this paper, we present a practical approach to adaptive depth networks that is applicable to both convolutional neural networks (CNNs) and transformers with minimal training effort.
 In our approach, every hierarchical residual stage is divided into two sub-paths, and they are trained to acquire different properties through a simple self-distillation strategy.
-While the first sub-path is essential for hierarchical feature learning, the second one is trained to refine the learned features and minimize performance degradation if it is skipped.
-Unlike prior adaptive networks, our approach does not train every target sub-network in an iterative manner.
+While the first sub-path is essential for hierarchical feature learning, the second one is trained to refine the learned features and minimize performance degradation even if it is skipped.
+Unlike prior adaptive networks, our approach does not train every target sub-network exhaustively.
 At test time, however, we can connect these sub-paths in a combinatorial manner to select sub-networks of various accuracy-efficiency trade-offs from a single network.
 We provide a formal rationale for why the proposed training method can reduce overall prediction errors while minimizing the impact of skipping sub-paths.
-We demonstrate the generality and effectiveness of our approach with convolutional neural networks and transformers.
+We demonstrate the generality and effectiveness of our approach with both CNNs and transformers.
+Source codes are available at https://github.com/wchkang/depth
 1 Introduction
-Modern deep neural networks such as convolutional neural networks (CNNs) and transformers [1] provide state-of-the-art performance at high computational costs, and, hence, lots of efforts have been made to leverage those inference capabilities in various resource-constrained devices.
+Modern deep neural networks such as CNNs and transformers [1] provide state-of-the-art performance at high computational costs, and, hence, lots of efforts have been made to leverage those inference capabilities in various resource-constrained devices.
 Those efforts include compact architectures [2, 3], network pruning [4, 5], weight/activation quantization [6], knowledge distillation [7], to name a few.
@@
 Moreover, unlike width-adaptation networks, there is no established principle for adapting depths since the impact of skipping individual layers has not been formally defined.
-Figure 1: (a) During training, each residual stage of a network is divided into two sub-paths.
-The layers in the second sub-paths are optimized to minimize performance degradation even if they are skipped.
-(b) At test time, we can instantly select $2^{N_{r}}$ parameter-sharing sub-networks from a single network by systematically skipping the last half layers of $N_{r}$ residual stages.
-(c) The sub-networks selected from a single network form a better Pareto frontier compared to individual networks.
+Figure 1: (a) During training, every residual stage of a network is divided into two sub-paths.
+The layers in every second (orange) sub-path are optimized to minimize performance degradation even if they are skipped.
+(b) At test time, these second sub-paths can be skipped in a combinatorial manner, allowing instant selection of various parameter sharing sub-networks.
+(c) The sub-networks selected from a single network form a better Pareto frontier than counterpart individual networks.
 In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to various networks, e.g., CNNs and transformers.
@@
 The proposed self-distillation strategy does not require exhaustive training of every target sub-network, resulting in significantly shorter training time than prior adaptive networks.
-Nevertheless, at test time, sub-networks with various depths can be selected instantly from a single network by connecting these sub-paths in a combinatorial manner, as shown in Figure 1-(b).
+However, at test time, sub-networks with various depths can be selected instantly from a single network by connecting these sub-paths in a combinatorial manner, as shown in Figure 1-(b).
 Further, these sub-networks with varying depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
@@
 In Section 4, we empirically demonstrate that our adaptive depth networks outperform counterpart individual networks, both in CNNs and vision transformers, and achieve actual inference acceleration and energy-saving.
-To authors’ best knowledge, this is the first approach to adaptive networks that provides a general principle and theoretic basis for how depth adaptation can be supported with minimal performance degradation and training effort.
-This allows us to avoid typical exhaustive training of target sub-networks and instead construct sub-networks of varying depths from specifically trained sub-paths.
+Our approach uniquely introduces a principle for training selected sub-paths to be skippable with minimal performance loss.
+This principle allows us to avoid typical exhaustive training of target sub-networks and instead construct sub-networks of varying depths from specifically trained sub-paths.
 We anticipate these advances will facilitate future research and applications of adaptive networks.
@@
 Dynamic Networks: Dynamic networks [21] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaptation of CNN models [22, 23, 24, 25, 26] and transformers [27, 28, 29, 30].
-In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layers (or blocks) that determine if the layers can be skipped [31, 22, 32, 33].
+In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layer (or block) that determine if the layer can be skipped [31, 22, 32, 33].
 These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs.
 However, the learned policy for skipping layers is opaque to users and does not provide a formal description of which layers can be skipped for a given input.
-Therefore, the network depths cannot be adapted in a predictable manner to meet the resource condition of target devices.
+Therefore, the network depth cannot be controlled by users in a predictable manner to meet the resource condition of target devices.
 Residual Blocks with Shortcuts: Since the introduction of ResNets [34], residual blocks with identity shortcuts have received extensive attention because of their ability to train very deep networks, and have been chosen by many CNNs [35, 36] and transformers [1, 37, 38].
@@
 The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using the proposed self-distillation strategy.
-Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input $\mathbf{h}^{s}$ to the next residual stage.
-In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped.
-This architectural pattern is applicable to both CNNs and transformers.
+Having similar distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as input $\mathbf{h}^{s}$ to the next residual stage.
+In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited if the second sub-path is skipped.
+These sub-paths are building blocks to construct sub-networks of varying depths.
 3.1 Architectural Pattern for Depth Adaptation
-In hierarchical networks, such as ResNets [34] and Swin transformers [38], there are typically 4 to 7 consecutive residual stages.
-In a network with $N_{r}$ residual stages, the $s$-th ($s=1,...,N_{r}$) residual stage consists of $L$ identical residual blocks  [^1] In transformers, residual blocks are called encoder-decoder blocks.
-In most networks, $L$ is usually 3 to 6.
-that transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
+In modern hierarchical networks, such as ResNets [34] and Swin transformers [38], there are typically 4 to 7 consecutive residual stages.
+[^1] For vision transformers, residual blocks/stages refer to encoder blocks/stages.
+In a network with $N_{r}$ stages, the $s$-th ($s=1,...,N_{r}$) stage consists of $L$ (typically $L\in[3,6]$) identical residual blocks that transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
 $\underbrace{\mathbf{h}^{s}_{1}}_{=\mathbf{h}^{s-1}}+\underbrace{F_{1}(\mathbf{h}^{s}_{1})+...+{F}^{s}_{L/2}(\mathbf{h}^{s}_{L/2})}_{\mathbf{F}^{s}_{base}}+\underbrace{F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})...+F_{L}(\mathbf{h}^{s}_{L})}_{\mathbf{F}^{s}_{skippable}}=\mathbf{h}^{s}$ (1)
@@
 To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2.
-We train these two sub-paths to have different properties.
+We train these two sub-paths to have different properties (Section 3.2).
 While $\mathbf{F}^{s}_{base}$ is trained to learn feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}(=\mathbf{h}^{s}_{L+1}$).
 Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference.
-In contrast, layers in $\mathbf{F}^{s}_{skippable}$ can be skipped to save computation since they only refine $\mathbf{h}^{s}_{base}$.
-If $\mathbf{F}^{s}_{skippable}$ is skipped, then intermediate features $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage.
-Therefore, $2^{N_{r}}$ sub-networks with different accuracy-efficiency trade-offs can be selected from a single network by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{r})$ or not.
+In contrast, layers in $\mathbf{F}^{s}_{skippable}$ can be skipped for efficiency since they only refine $\mathbf{h}^{s}_{base}$.
+If $\mathbf{F}^{s}_{skippable}$ is skipped, then $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage.
+Therefore, $2^{N_{r}}$ (=$\sum_{k=0}^{N_{r}}{C(N_{r},k)}$) sub-networks with varying accuracy-efficiency trade-offs can be selected from a single network by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{r})$ or not (Table 5).
+This architectural pattern is agnostic to the type of residual blocks.
+The residual function $F$ can be convolution layers for CNNs and self-attention + MLP layers for transformers.
 Table 4 in Appendix A.1 provides details on how this pattern is applied to both CNNs and transformers.
@@
 In Algorithm 1, the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are exploited.
-During the forward passes of the super-net and the base-net in steps 5 and 8, respectively, intermediate features $\mathbf{h}_{super}$ and $\mathbf{h}_{base}$ are obtained.
-In steps 9-10, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in the loss function is explicitly minimized for training data $\mathbf{X}$.
-Further, this self-distillation step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage.
-Therefore, $\mathbf{h}^{s}_{base}$ is expected to learn more compact representation from $\mathbf{h}^{s}_{super}$.
-Self-distillation itself has been extensively used in prior adaptive networks [11, 10, 13].
+In steps 9-10, the hierarchical features (and outputs) from the super-net and the base-net are encouraged to have similar distributions through self-distillation.
+Due to the architectural pattern of interleaving the mandatory and the skippable sub-paths, minimizing $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ also minimizes $D_{KL}(\mathbf{h}_{super}\|\mathbf{h}_{base})$ implicitly.
+Therefore, only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ can be included in $loss_{base}$ if the extraction of intermediate features is tricky.
+Self-distillation has been extensively used in prior adaptive networks [11, 10, 13].
 However, their primary goal is to train every target sub-network.
-For example, on each iteration, sub-networks are randomly sampled from a large search space to act as either teachers or students [11], which results in significantly longer training time.
+For example, on each iteration, sub-networks are randomly sampled from a large search space to act as either teachers or students [11], which takes significantly longer training time and can generate conflicting training objectives for different parameter-sharing sub-networks [15, 16].
 In contrast, our method in Algorithm 1 focuses on training sub-paths by exclusively using the super-net as a teacher and the base-net as a student.
-By focusing on training sub-paths rather than target sub-networks, the training procedure in Algorithm 1 is significantly simplified.
-At test time, however, $2^{N_{r}}(=\sum_{k=0}^{N_{r}}{C(N_{r},k)})$ different sub-networks can be constructed instantly by connecting these sub-paths in a combinatorial manner.
+By focusing on sub-paths, the training procedure in Algorithm 1 is significantly simplified and avoids potential conflicts among sub-networks.
+At test time, however, various sub-networks can be selected by connecting these sub-paths.
 The effect of this self-distillation strategy is investigated in Section 4.4.
@@
 In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude.
-In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the proposed self-distillation strategy, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation; see Figure 3 for empirical evidence.
+In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the proposed self-distillation strategy (refer to Figure 3 for empirical evidence).
+As a result, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation.
 If we similarly keep expanding the loss function around $\mathbf{h}^{s}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
@@
 Considering this result, we can conjecture that, with our architectural pattern and self-distillation strategy, layers in $\mathbf{F}^{s}_{skippable}$ learn functions that refine input features $\mathbf{h}^{s}_{base}$ iteratively for better inference accuracy while minimally changing the distribution of $\mathbf{h}^{s}_{base}$.
-Figure 3: $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at residual blocks. In ours, skippable sub-paths (orange areas) minimally change the distribution of input h
+Figure 3: $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at residual blocks. In ours, skippable sub-paths (orange areas) minimally change the distribution of input h.
 Empirical Analysis: We can estimate how much the distribution of input $\mathbf{h}$ is transformed by residual function $F$ by measuring $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at each residual block.
-Figure 3 illustrates $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at every residual block in the baseline ResNet50 and our ResNet50-ADN that is trained according to Algorithm 1  [^2] ImageNet validation dataset is used for this experiment..
+Figure 3 illustrates $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at every residual block in the baseline ResNet50 and our ResNet50-ADN that is trained according to Algorithm 1.222ImageNet validation dataset is used for this experiment.
 Our ResNet50-ADN exhibits greater transformation than ResNet50 in the mandatory sub-paths (blue areas) and less transformation in the skippable sub-paths (orange areas).
@@
 4 Experiments
-We use four representative residual networks both from CNNs and vision transformers as base models to apply the proposed architecture pattern: MobileNet V2 [35] is a lightweight CNN model, ResNet [34] is a larger CNN model, and ViT [37] and Swin-T [38] are representative vision transformers.
+We use networks both from CNNs and vision transformers as base models to apply the proposed architecture pattern: MobileNet V2 [35] is a lightweight CNN model, ResNet [34] is a larger CNN model, and ViT [37] and Swin-T [38] are representative vision transformers.
 All base models except ViT have hierarchical stages, each with 2 $\sim$ 6 residual blocks.
-So, according to the proposed architectural pattern, every residual stage is evenly divided into 2 sub-paths for depth adaptation.
+So, according to the proposed architectural pattern, every stage is evenly divided into 2 sub-paths for depth adaptation.
 ViT does not define hierarchical stages and all residual encoder blocks have same spatial dimensions.
@@
 Details are in Table 4 in Appendix A.1.
+These models are trained according to Algorithm 1.
+For self-distillation, only the final outputs from the super-net and the base-net, or $\mathbf{\hat{y}}_{super}$ and $\mathbf{\hat{y}}_{base}$ respectively, are used since exploiting intermediate features for explicit self-distillation has a marginal impact on performance (Section 4.4).
 We use the suffix ‘-ADN’ to denote our adaptive depth networks.
-Since our adaptive depth networks have many parameter-sharing sub-networks in a single network, we indicate which sub-network is used for evaluation in parenthesis, e.g., ResNet50-ADN (super-net).
-For our sub-networks, boolean values are also used to indicate in which residual stages the sub-paths are skipped.
-For example, ResNet50-ADN (base-net) is equivalent to ResNet50-ADN (TTTT).
+A series of boolean values in parentheses denotes a specific sub-network used for evaluation; each boolean value represents the residual stage where its skippable sub-path is skipped.
+For example, ResNet50-ADN (FFFF) and ResNet50-ADN (TTTT) correspond to the super-net and the base-net of ResNet50-ADN, respectively.
 Model Params FLOPs Acc
@@
 ResNet50-Base 17.11 2.58 75.0
-MbV2-ADN (FFFF) 3.72 0.32 72.7
-MbV2-ADN (TTTT) 0.25 70.7
+MbV2-ADN (FFFFF) 3.53 0.32 72.5
+MbV2-ADN (TTTTT) 0.22 70.6
 MbV2 3.50 0.32 72.1
-MbV2-Base 2.99 0.25 70.7
-ViT-b/16-ADN (FFFF) 86.59 17.58 81.3
-ViT-b/16-ADN (TTTT) 11.76 79.4
+MbV2-Base 2.98 0.22 70.2
+ViT-b/16-ADN (FFFF) 86.59 17.58 81.4
+ViT-b/16-ADN (TTTT) 11.76 80.6
 ViT-b/16 86.57 17.58 81.1
 ViT-b/16-Base 67.70 11.76 78.7
-Swin-T-ADN (FFFF) 28.30 4.49 81.8
-Swin-T-ADN (TTTT) 2.69 78.7
+Swin-T-ADN (FFFF) 28.30 4.49 81.6
+Swin-T-ADN (TTTT) 2.34 78.0
 Swin-T 28.29 4.49 81.5
-Swin-T-Base 15.63 2.69 78.6
+Swin-T-Base 15.34 2.34 77.4
 Figure 4: (a) Results on ImageNet validation dataset.
@@
 In Figure 4-(b) and Table 1, several state-of-the-art efficient inference methods and dynamic networks are compared with our base-networks.
-The result demonstrates that our adaptive depth networks show competitive performance across a range of FLOPs.
+The result demonstrates that our adaptive depth networks show comparable performance across a range of FLOPs.
 In Figure 4-(b), it should be noted that individual ResNets trained with knowledge distillation has worse performance than individual ResNets.
@@
 AlphaNet-0.75x [11] $\ast$ 0.21G 34% 70.5%
-MbV2-ADN (TTTT) 0.25G 22% 70.7%
+MbV2-ADN (TTTTT) 0.22G 32% 70.6%
 WDPruning-0.3-11 [57$\dagger$ 2.6G - 78.3%
@@
 SPViT-Swin-Ti [59$\star$ 3.3G 27% 80.1%
-Swin-T-ADN (TTTT) 2.7G 40% 78.7%
-4.2 Training Time
-One major advantage of our adaptive depth networks is that they require significantly shorter training time than prior adaptive networks.
-Figure 5-(a) shows that training our adaptive depth networks take similar time to train two individual networks combined.
-This is because our training method in Algorithm 1 exclusively exploits only the super-net and the base-net for self-distillation.
-In contrast, compared adaptive networks require much longer training time than ours since they have to train every target sub-network explicitly  [^3] Although the prior works we compared for training time may seem outdated, they remain relevant and representative because there has been little progress in reducing the training time of adaptive networks..
+Swin-T-ADN (TTTT) 2.3G 48% 78.0%
+4.2 Training Time and Sub-Networks
+One of the key advantages of our training method in Algorithm 1 is its simplicity and significantly reduced training effort compared to previous adaptive networks.
+Figure 5-(a) shows that the training time of our adaptive depth networks are similar to the training time of two individual networks combined.
+This is because our training method trains sub-paths, rather than sub-networks, by exploiting only the super-net and the base-net.
+In contrast, the compared adaptive networks require much longer training time than ours since they have to explicitly apply self-distillation to all target sub-networks.
 For example, on every training iteration, AlphaNet [11] randomly samples sub-networks from its search space for self-distillation.
+Although the prior works we compared may seem outdated, they remain relevant and representative because there has been little progress in improving the training cost of adaptive networks.
 While AlphaNet [11] supports significantly more sub-networks, it should be noted that supporting many sub-networks is not our goal.
 Instead, our objective is to provide better performance Pareto with several useful sub-networks, as demonstrated in Figure 4-(b).
-For example, although ResNet50-ADN can select $2^{4}$ sub-networks, Table 5 (in Appendix A.2) shows that some sub-networks are more useful than the others; results in Table 5 show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
+For example, although ResNet50-ADN has $2^{4}$ sub-networks in its search space, Table 5 in Appendix A.2 demonstrates that some sub-networks manifest better performance than the others even if they have similar depths.
+The results show that skipping sub-paths in deeper residual stages generally has a more detrimental impact on performance, suggesting that feature refinement is more crucial for deeper residual stages.
 Model 1 epoch (min) # of sub-nets
@@
 MbV2-Base 10.0 -
-MbV2-ADN (ours) 22.5 $2^{4}$ depths
+MbV2-ADN (ours) 22.5 $2^{5}$ depths
 AlphaNet [11] 230.5 216 depths
@@
 While the Pareto frontiers formed by sub-networks demonstrate theoretic performance, inference acceleration in actual devices is more important in practice for effective control of inference latency and energy consumption.
-Figure 5(b) shows the performance in Nvidia Jetson Orin Nano.
+Figure 5-(b) shows the performance on Nvidia Jetson Orin Nano.
 The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50, a representative width-adaptation network.
 The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption.
-For example, in ResNet50-ADN, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%.
+Although our ResNet50-ADN has a limited FLOPs adaptation range, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%.
 In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
 4.4 Ablation Study
-Table 2: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. $\downarrow$ and $\uparrow$ in parentheses are comparisons to non-adaptive individual networks.
-self- skip-aware ResNet50 Acc@1 (%) ViT-b/32 Acc@1 (%)
+Table 2: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN.
+Applied components are checked.
+$\downarrow$ and $\uparrow$ in parentheses are comparisons to non-adaptive individual networks.
+By default, only the outputs, or $\mathbf{\hat{y}}_{super}$ and $\mathbf{\hat{y}}_{base}$, are used for self-distillation.
+The last row with double check marks shows the results when both intermediate features and outputs are used for self-distillation.
+self- skip-aware ResNet50-ADN Acc@1 (%) ViT-b/32-ADN Acc@1 (%)
 distllation. BNs/LNs FFFF TTTT FFFF TTTT
 75.2% ($\downarrow$ 1.5%) 72.2% ($\downarrow$ 2.8%) 75.7% ($\downarrow$0.2%) 74.1% ($\uparrow$ 0.3%)
-O 76.1% ($\downarrow$ 0.6%) 74.9 % ($\downarrow$ 0.1%) 76.4% ($\uparrow$0.5%) 74.3% ($\uparrow$0.5%)
-O 76.6% ($\downarrow$ 0.1%) 75.1% ($\uparrow$ 0.1%) 76.0% ($\uparrow$0.1%) 74.3% ($\uparrow$0.5%)
-O O 77.6% ($\uparrow$ 0.9%) 76.1% ($\uparrow$ 1.1%) 76.6% ($\uparrow$ 0.8%) 74.3% ($\uparrow$0.5%)
+76.1% ($\downarrow$ 0.6%) 74.9 % ($\downarrow$ 0.1%) 76.4% ($\uparrow$0.5%) 74.3% ($\uparrow$0.5%)
+76.6% ($\downarrow$ 0.1%) 75.1% ($\uparrow$ 0.1%) 76.0% ($\uparrow$0.1%) 74.3% ($\uparrow$0.5%)
+77.6% ($\uparrow$ 0.9%) 76.1% ($\uparrow$ 1.1%) 76.6% ($\uparrow$ 0.8%) 74.3% ($\uparrow$0.5%)
+✓✓ 77.3% ($\uparrow$ 0.6%) 76.2% ($\uparrow$ 1.2%)
 We first investigate the influence of two key components of the proposed adaptive depth networks: (1) self-distillation of sub-paths and (2) skip-aware BNs/LNs.
@@
 Table 2 shows the results.
-For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than non-adaptive individual networks by 1.5% and 2.8%, respectively.
+For ResNet50-ADN, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than non-adaptive individual networks by 1.5% and 2.8%, respectively.
 This result shows the difficulty of joint training sub-networks for adaptive networks.
 When one of the two components is applied individually, the performance is still slightly worse than individual networks’.
-In the last row, when both self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net.
-The results with ViT-b/32 show that switchable layer normalization has similar effect in vision transformers.
+When both self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net.
+The last row, with double check marks, shows that exploiting intermediate features as well as softmax outputs for self-distillation has minor impact on performance.
 Table 3: Comparison of self-distillation strategies. Our approach (in bold) uses exclusively the super-net and the base-net as a teacher and a student, respectively.
-sub-nets ResNet50-ADN Acc@1 (%) ViT/b-32-ADN Acc@1 (%)
+sub-nets ResNet50-ADN Acc@1 (%) ViT-b/32-ADN Acc@1 (%)
 Teacher Student FFFF TFFF TTFF TTTF TTTT FFFF TFFF TTFF TTTF TTTT
-FFFF TTTT 77.6 77.2 76.8 76.1 76.1 76.6 76.0 75.5 74.5 74.3
+FFFF TTTT 77.6 77.1 76.5 76.0 76.1 76.6 76.0 75.5 74.5 74.3
 FFFF Random 77.1 76.7 76.4 75.5 74.8 75.2 74.7 73.9 72.9 71.1
@@
 These randomly sampled sub-networks are trained explicitly through self-distillation.
-However, as shown in Table 3, all sub-networks, both from ResNet50-ADN and ViT/b-32-ADN, perform significantly better when our self-distillation strategy is applied for training.
-Even though some sub-networks (such as TTTF, TTFF, and TTTF) are not trained explicitly with our self-distillation strategy, they still outperform their counterpart sub-networks trained through random sampling.
+However, as shown in Table 3, all sub-networks, both from ResNet50-ADN and ViT-b/32-ADN, perform significantly better when our self-distillation strategy is applied for training.
+Even though most of our sub-networks (such as TFFF, TTFF, and TTTF) are instantly constructed at test time without explicit training, they still outperform their counterpart sub-networks trained explicitly through random sampling.
 This result demonstrates that our method of training sub-paths is more effective than training target sub-networks themselves.
-Figure 6: Pareto frontiers of ResNet50-ADNs trained with varying ratios between mandatory and skippable sub-paths. Total number of blocks remains unchanged.
-Lengths of Mandatory Sub-Paths: In Figure 6, we investigate the effect of varying the length ratio between the mandatory and the skippable sub-paths.
+Figure 6: Pareto frontiers of three ResNet50-ADNs, each trained with varying ratios between mandatory and skippable sub-paths. Total number of blocks remains unchanged.
+Lengths of Mandatory Sub-Paths: In Figure 6, we investigate the impact of varying the ratio of lengths between the mandatory and the skippable sub-paths.
 (Details are in Table 6.) If mandatory sub-paths become shorter than skippable sub-paths, smaller sub-networks can be selected since more layers can be skipped.
-However, the result in Figure 6 shows that this configuration significantly degrades performance of all sub-networks.
+However, the result in Figure 6 shows that this configuration (shown in blue line) significantly degrades performance of all sub-networks.
 Since every sub-network shares parameters of mandatory sub-paths, low inference capability of shallow mandatory sub-paths affects all sub-networks.
-This implies that maintaining certain depths in mandatory sub-paths is essential for effective inference.
-In contrast, increasing the length of mandatory sub-paths further does not yield further performance improvement and only reduce the range of FLOPs adaptation.
+This implies that maintaining certain depths in mandatory sub-paths is crucial for effective inference.
+Conversely, further increasing the length of mandatory sub-paths (shown in green line) does not further improve performance and only reduces the range of depth adaptation.
 5 Conclusions
 We propose a practical approach to adaptive depth networks that can be applied to both CNNs and transformers with minimal training effort.
-We provide a general principle and formal explanation on how depth adaptation can be achieved with minimal performance degradation.
-Under this principle, our approach can avoid typical exhaustive training of target sub-network and instead focus on optimizing the sub-paths of the network to have specific properties.
-At test time, those sub-paths can be connected in a combinatorial manner to construct sub-networks with various accuracy-efficiency trade-offs from a single network.
-Experimental results show that those sub-networks form a better Pareto frontier than non-adaptive baseline networks and achieve actual inference acceleration.
+We provide a general principle and a formal explanation on how depth adaptation can be achieved with minimal performance degradation.
+Under this principle, our approach can avoid typical exhaustive training of target sub-networks and instead focus on optimizing the sub-paths of the network to have specific properties.
+At test time, these sub-paths can be connected in a combinatorial manner to construct sub-networks with various accuracy-efficiency trade-offs from a single network.
+Experimental results show that these sub-networks form a better Pareto frontier than non-adaptive baseline networks and achieve actual inference acceleration.
 Appendix A Appendix: Detailed Settings and Evaluation Results
@@
 stage 1 1 - 1
-stage 2 2 - 2
+stage 2 1 1 2
 stage 3 2 1 3
@@
 stage 3 3 3 6
-stage 4 2 - 2
+stage 4 1 1 2
 (d) Swin-T-ADN
-Table 4: Each residual stage of base models is evenly divided into two sub-paths; the first is mandatory and the other is skippable.
+Table 4: Each stage of base models is evenly divided into two sub-paths; the first is mandatory and the other is skippable.
 Since ViT does not define hierarchical stages, 12 identical encoder blocks are divided into 4 stages, resembling other residual networks for vision tasks.
@@
 FFFF 4.11 77.6
-TFFF 3.90 77.2
+TFFF 3.90 77.1
 FTFF 3.68 76.5
 FFTF 3.46 75.6
-FFFT 3.90 76.2
-TTFF 3.46 76.8
-TFTF 3.24 75.4
-TFFT 3.68 76.1
-FTTF 3.02 75.3
-FTFT 3.46 75.4
+FFFT 3.90 76.7
+TTFF 3.46 76.5
+TFTF 3.24 75.3
+TFFT 3.68 76.4
+FTTF 3.02 75.8
+FTFT 3.46 75.9
 FFTT 3.25 75.6
-TTTF 2.80 76.1
-TTFT 3.24 75.9
-TFTT 3.02 75.4
-FTTT 2.80 75.6
+TTTF 2.80 75.9
+TTFT 3.24 75.8
+TFTT 3.02 75.3
+FTTT 2.80 76.0
 TTTT 2.58 76.1
@@
 sub-network FLOPs (G) Acc@1 (%)
-FFFF 17.58 81.3
-TFFF 16.20 80.5
-FTFF 16.20 80.2
-FFTF 16.20 80.0
-FFFT 16.20 80.4
-TTFF 14.67 79.9
-TFTF 14.67 79.6
-TFFT 14.67 79.8
-FTTF 14.67 79.4
-FTFT 14.67 79.7
-FFTT 14.67 79.4
-TTTF 13.21 79.3
-TTFT 13.21 79.4
-TFTT 13.21 79.2
-FTTT 13.21 79.1
-TTTT 11.76 79.1
+FFFF 17.58 81.4
+TFFF 16.20 81.1
+FTFF 16.20 81.0
+FFTF 16.20 80.6
+FFFT 16.20 81.2
+TTFF 14.67 80.9
+TFTF 14.67 80.4
+TFFT 14.67 80.9
+FTTF 14.67 80.5
+FTFT 14.67 80.9
+FFTT 14.67 80.6
+TTTF 13.21 80.5
+TTFT 13.21 80.7
+TFTT 13.21 80.5
+FTTT 13.21 80.6
+TTTT 11.76 80.6
 (b) ViT-b/16-ADN
 Table 5: FLOPs and ImageNet validation accuracy of sub-networks.
-Only the super-net and the base-net are trained explicitly.
-The other parameter-sharing sub-networks in the middle can be selected at test time without explicit training.
+Only super-net (or, FFFF) and base-net (or, TTTT) are trained explicitly.
+Sub-networks in the middle can be selected at test time without explicit training.
 The highest accuracy in each group is shown in bold.
-This results show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
-A.3 Varying the Lengths of Sub-Paths in ResNet50-ADN
+A.3 Varying the Ratio of Sub-Path Lengths in ResNet50-ADN
 number of number of total
@@
 (b) # of mandatory $>$ # of skippable
-Table 6: Varying the length ratio of sub-paths of ResNet50-ADN. Total number of blocks of each stage is not changed.
-A.4 Varying Sub-Paths Length Ratios in Vit-b/16-ADN
+Table 6: The configurations of ResNet50-ADNs with different proportions between mandatory and skippable sub-paths. Total number of blocks at each stage remains unchanged.
+A.4 Varying the Ratio of Sub-Path Lengths in Vit-b/16-ADN
 Each stage of Vit-b/16-ADN has 3 encoder blocks and, by default, we select every last encoder block as a skippable sub-path.
@@
 As demonstrated with ResNet50-ADN in Figure 6, low inference capability of shallow mandatory sub-paths affects all sub-networks.
-This again shows that some minimum depths in mandatory sub-paths is required for effective inference.
+This result again shows that maintaining certain depths in mandatory sub-paths is crucial for effective inference.
 number of number of total
@@
 [61] ResNet50-Base 228.73G 32.7 29.9 34.1 31.2
+RetinaNet ResNet50 151.54G 36.4 37.4
+[62] ResNet50-Base 132.04G 31.7 35.2
 In order to investigate the generalization ability of our approach, we use MS COCO 2017 datasets on object detection and instance segmentation tasks using representative detectors.
@@
 (a) Original ResNet50
-(b) ResNet50-ADN (FFFF), or super-net
-(c) ResNet50-ADN (TTTT), or base-net
+(b) ResNet50-ADN (FFFF)
+(c) ResNet50-ADN (TTTT)
 Figure 8: Class Activation Maps of the 3rd residual stages of ResNet50s.
 (a) Original ResNet50’s activation regions change gradually across all blocks.
-(b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features.
+(b) In ResNet50-ADN (FFFF), the first 3 blocks have extensive hot activation regions, implying active learning of new level features.
 In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target.